### v1

# CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN

Abstract
As robotic arm applications extend beyond industrial settings into healthcare, service, and daily life, existing control algorithms struggle to achieve the agile manipulation required for complex environments with dynamic trajectories, unpredictable interactions, and diverse objects. This paper presents a biomimetic control framework based on Spiking Neural Networks (SNN), inspired by the human Central Nervous System (CNS), to achieve agile control in such environments. The proposed framework features five control modules (cerebral cortex, cerebellum, thalamus, brainstem, spinal cord), three hierarchical control levels (first-order, second-order and third-order), and two information pathways (ascending and descending). Each module is fully implemented using SNN. The spinal cord module uses spike encoding and Leaky Integrate-and-Fire (LIF) neurons for feedback control. The brainstem module employs a network of LIF and non-spiking LIF neurons to dynamically adjust spinal cord parameters via reinforcement learning. The thalamus module similarly adjusts the cerebellum’s torque outputs. The cerebellum module uses a recurrent SNN to learn the robotic arm’s dynamics through regression, providing feedforward gravity compensation torques. The framework is validated both in simulation and on real-world robotic arm platform under various loads and trajectories. Results demonstrate that our method outperforms the industrial-grade position control in manipulation agility.
I Introduction
With the advancement of actuator and material technologies, robotic arms have gradually increased in degrees of freedom while decreasing in cost, enabling them to perform more delicate operations. Consequently, their applications have expanded in recent years to fields such as healthcare, services, and entertainment. As robotic arms become increasingly integrated into our daily lives, the complexity of their operating environments and the agility required for diverse, fine-grained tasks pose new challenges to control strategies. For instance, a 7-DOF robotic arm can readily cut a circular contour on a workpiece or transport heavy objects, yet it still struggles to perform a task as subtle as peeling an apple—a task easily accomplished by the human arm. Compared with biological systems, even the most advanced robotic arms today fall short of the agility of the human upper limb [mason2018], underscoring the existence of advanced motor control mechanisms in humans that remain underutilized in robotic systems.
In well-modeled environments, control algorithms combining Whole-Body Control (WBC) [Abe2007, liu2011interactive, weighttans2011, Lee2012Intermediate, Escande2014, Kim2019, han2021recursive] and Model Predictive Control (MPC) [carron2019data, pankert2020perceptive, wang2023trajectory] can effectively counteract disturbances and achieve strong control performance. Such model-based approaches reflect human capabilities in applying prior knowledge to solve specific tasks. However, they lack the adaptability inherent to human motor learning. In complex scenarios, variations in the size, mass, or dynamics of manipulated objects may render the kinematic and dynamic models inaccurate, limiting the effectiveness of purely model-based strategies.
With the rapid progress of neural networks, their powerful approximation and learning capabilities have motivated researchers to employ them for adaptive model correction in response to environmental changes. For example, [chaoui2009ann] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [he2017adaptive] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [wang2018adaptive] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [salloom2020adaptive] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [liu2019adaptive] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [pham2020adaptive] applied radial basis function networks with dynamic surface control for robust dual-arm coordination. While these approaches achieved progress, they rely on artificial neural networks (ANN), which inherently disregard the temporal-spatial characteristics of motion, focusing only on instantaneous control. Furthermore, they remain constrained by robotic dynamics models, with network structures and training methods often tailored to specific problems and tightly coupled to complex mathematical formulations, thereby limiting generalizability.
Recently, spiking neural networks (SNN) combined with brain-inspired architectures have emerged as promising alternatives. By emulating biological neurons, SNNs process and store information through spike trains and membrane potential dynamics, offering advantages in computational speed, energy efficiency, information dimension, and biological plausibility, making them well-suited for robotic control [bing2018survey]. For instance, [carrillo2008real] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn manipulator dynamics for torque control; [abadia2019robot] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [abadia2021cerebellar] further established resilience to control delays, highlighting potential in teleoperation and cloud robotics and refines this framework to enable robotic arms to dynamicallyy adjust compliance. However, these methods suffer from high structural complexity, computational cost, hyperparameter sensitivity, and limited interpretability, as they focus narrowly on cerebellar functionality while neglecting the integrative role of the broader central nervous system (CNS), which coordinates control through interactions among several neural areas[Halassa_2022].
Inspired by this systemic perspective of the CNS, [li2023cbmc] proposed the CBMC framework, incorporating modules such as the cortex, cerebellum, and spinal cord. Yet its computational inefficiency restricted real-time applicability. Building on this, [10769842] introduced a hierarchical control architecture deployable on neuromorphic hardware, though its lack of neural networks in some modules compromised accuracy.
Despite these advances, current brain-inspired algorithms have yet to demonstrate agile performance comparable to that of humans. As defined in [8610238], agility is a capability characterized by rapid and efficient transitions between diverse motion patterns, with precision and speed as its key attributes. In this work, we propose CBMC-V3, a fully SNN-based brain-inspired control framework designed to achieve agile robotic arm control. The key contributions are fourfold: (1) Proposing a CNS-inspired control framework leveraging hierarchical modularity; (2) Designing a fully SNN-based implementation; (3) Validating the framework and module design in simulation experiments; (4) Deploying the algorithm on a robotic platform, demonstrating agility.
The remainder of this paper is organized as follows: Section II introduces the framework structure; Section III details the SNN-based implementations of individual modules; Section IV presents simulation analyses; Section V conducts experiments on a robotic arm platform, and Section VI concludes the study.
Figure 1: Overview of the human CNS and the proposed framework structure. (a) Human CNS relating to motion control. Sensory information and motor commands are transmitted through three orders of neurons along ascending and descending pathways. Sensory signals travel from peripheral receptors to the spinal cord, brainstem, and ultimately the cerebellum and cerebral cortex, while motor commands descend from the cerebral cortex and cerebellum through the thalamus and spinal cord to the muscles. The spinal cord can independently generate reflexive responses via reflex arcs, enabling rapid reactions to stimuli [LLOBERA2023190]. The brainstem contributes to movement termination, reorientation, and coordination with the spinal cord [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137]. The thalamus functions as a hub for motor coordination, multimodal perception, and sensorimotor integration [LLOBERA2023190, Halassa_2022]. The cerebellum regulates body movements, balance, and posture by integrating sensory inputs and refining motor commands [LLOBERA2023190]. The cerebral cortex generates decisions and motion trajectories. (b) Schematic diagram of our proposed framework, featuring five modules, three control levels and two pathways as mentioned in (a). (c) Neural network design of each module, utilizing three types of spiking neurons: input fiber (grey), LIF neuron (yellow) and non-spiking LIF neuron (green).
II Framework structure
Inspired by the motion control mechanism of the human CNS, as illustrated in Fig.1a, the structure of our proposed framework CBMC-V3 is shown in Fig.1b. This framework consists of three fundamental components: (1) five functional modules, namely the cerebral cortex module, cerebellum module, thalamus module, brainstem module, and spinal cord module; (2) three hierarchical control levels, namely the first-order, second-order, and third-order level; (3) two information pathways, the ascending and descending pathway. In the following sub-sections, we will introduce each of these components in detail and explain how the framework integrates functional modules, control loops, and information pathways to accomplish a complete operational cycle.
II-A Module function
In the human CNS, the spinal cord can generate reflexive actions independently via reflex arcs, enabling rapid responses to sensory stimuli without cortical input [LLOBERA2023190]. Inspired by this mechanism, the spinal cord module in our framework implements a PD-like feedback control function, producing torque commands at real-time control frequencies, i.e.,
$\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ $=\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})(\bm{\dot{q}_{d}}-\bm{\dot{q}})$ (1)
$+\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})(\bm{q_{d}}-\bm{q})$
The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137]. Motivated by this, the brainstem module in our framework dynamically adjusts the PD gains for each joint based on trajectory information,i.e.,
$\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{v,1},k_{v,2},\cdots,k_{v,n})$ (2)
$\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{p,1},k_{p,2},\cdots,k_{p,n})$ (3)
providing the spinal cord module with time-varying proportional and derivative coefficients, i.e., $k_{p,i}$ and $k_{v,i}$ for $i=1,\dots,n$. By integrating spinal feedback control with brainstem-driven adaptive gain modulation, the framework circumvents the complexity of manipulator dynamics while compensating for structured uncertainties through learning.
The gravity compensation torque of a robotic manipulator, denoted as $\bm{\tau}_{g}(\bm{q})$, is a feedforward torque designed to counteract the gravitational forces acting on the manipulator itself. It plays a crucial role in ensuring the smoothness and stability of motion. In human motor control, the cerebellum is primarily responsible for coordinating and regulating body movements, ensuring muscle and postural balance. It integrates sensory inputs, such as muscle tension and joint position, and adjusts the magnitude and direction of motor commands to achieve precise and fluid motion [LLOBERA2023190]. Analogously, gravity compensation in robotic manipulators serves a function comparable to that of the cerebellum, and thus the cerebellar module in our framework is tasked with providing this compensation. Furthermore, we note that during human arm manipulation, accurate models of dynamics are not strictly required. Instead, the brain relies on stored motor patterns to cope with different loading conditions. For instance, when grasping an object, we estimate its weight based on prior experience and adjust the exerted force accordingly. Such experiential knowledge is typically not acquired on site during execution, but rather accumulated over time through long-term daily interactions. Inspired by this biological mechanism, the proposed framework equips the cerebellum module with the ability to pre-learn multiple gravity compensation ‘patterns’ corresponding to different load masses. These patterns can then be retrieved and combined to generate an $n\times m$ set of compensation torques,
$\begin{split}\bm{T}_{{\text{cb}}}(\bm{q})&=(\bm{\tau}_{\text{cb},1}(\bm{q}),\bm{\tau}_{\text{cb},2}(\bm{q}),\cdots,\bm{\tau}_{\text{cb},m}(\bm{q}))\\ &=\begin{pmatrix}\tau_{\text{cb},11}&\cdots&\tau_{\text{cb},1m}\\ \vdots&\ddots&\vdots\\ \tau_{\text{cb},n1}&\cdots&\tau_{\text{cb},nm}\end{pmatrix}\end{split}$ (4)
as a function of joint positions, where $\tau_{\text{cb},ji}(\bm{q}),j=1,\cdots,n,i=1,\cdots,m$, denotes the torque provided for the $j$-th joint under the $i$-th compensation pattern.
Although suitable motor patterns can often be selected based on prior experience, in certain cases, such selections may be inaccurate due to factors such as visual deception. For instance, the mass of an opaque box cannot be reliably estimated visually, which may lead to insufficient force to move the load or, conversely, an excessive initial force causing large errors. Nevertheless, humans can quickly adjust their motor output upon physically contacting the object, using sensory feedback to correct their movement patterns. This phenomenon is closely associated with the function of the thalamus, which serves as a central hub for motor coordination and plays a critical role in multimodal perception and sensorimotor integration [LLOBERA2023190, Halassa_2022]. Analogously, a robotic arm can ‘perceive’ the load weight via force sensors installed at the end-effector. The thalamus module, similar to humans, uses the sensed load information to dynamically weight and integrate the torque contributions provided by the cerebellar module, generating an actual gravity compensation torque, i.e.,
$\begin{split}\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})&=\bm{T}_{\text{cb}}(\bm{q})\cdot\bm{w}(\bm{F}_{\text{ef}})\\ &=\begin{pmatrix}\tau_{g,1}\\ \vdots\\ \tau_{g,n}\end{pmatrix}=\begin{pmatrix}w_{1}\tau_{\text{cb},11}+\cdots+w_{m}\tau_{\text{cb},1m}\\ \vdots\\ w_{1}\tau_{\text{cb},n1}+\cdots+w_{m}\tau_{\text{cb},nm}\end{pmatrix}\end{split}$ (5)
where $\bm{F}_{\text{ef}}$ denotes the end-effector torque sensor feedback, $\bm{w}(\bm{F}_{\text{ef}})=(w_{1},w_{2},\dots,w_{m})^{T}$ represents the dynamically learned weight coefficients of the thalamus module, and $\tau_{g,1},\dots,\tau_{g,n}$ are the resulting gravity compensation torques. By combining the prior knowledge stored in the cerebellum module with the adaptive weighting capability of the thalamus module, the framework effectively addresses the unknown dynamics of the manipulator’s end-effector load.
The resulting torque is then summed with the feedback torque generated by the spinal cord module and applied as joint-level control commands to the actuators, i.e.,
$\bm{\tau}=\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})+\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{1},\cdots,\tau_{n})^{T}$ (6)
where $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ is produced by the spinal cord and brainstem module, $\bm{\tau}g(\bm{q},\bm{F}_{\text{ef}})$ is produced by the cerebellum and thalamus module, and $(\tau_{1},\dots,\tau_{n})^{T}$ represents the joint torques delivered to the actuators.
In conclusion, the function of each module can be summarized as follows:
  • The cerebral cortex module makes decisions and generates motion trajectories. In this research, it stores and replays pre-planned trajectories.
  • The cerebellum module is a cerebellar-like network pre-trained to store several fundamental gravity compensation motion patterns.
  • The thalamus module learns to generate compensation torque adaptable to varying loads by online RL using proprioception information and the cerebellum module output.
  • The brainstem module learns to provide appropriate feedback control coefficients by online RL using proprioception information.
  • The spinal cord module combines feedback control coefficients from the brainstem and torque from the thalamus module to generate real-time control torque commands.
II-B Hierarchical control level
Clearly, different neural regions in the central nervous system operate hierarchically during motion control, with each level specializing in particular tasks while communicating and cooperating through neural pathways. This hierarchical organization enables organisms to execute a wide range of motion tasks flexibly and stably across varying environments. Physiological studies indicate that the multilevel transmission of neural signals is a key component of this hierarchy. Proprioceptive information is conveyed from sensory receptors to the brain via a three-order neuronal pathway[william_ascending_pathway, LLOBERA2023190], as illustrated in Fig.1a. First-order neurons receive impulses from the skin and muscles and transmit them to the spinal cord. Some of this information is processed locally in the spinal cord and then relayed back to the muscles through these first-order neurons. The remaining signals pass through second-order neurons, which transmit impulses via the brainstem to the thalamus. Finally, third-order neurons carry these impulses to the cerebral cortex and cerebellum, and are responsible for information exchange between them.
Inspired by this three-order neuronal pathway, the proposed framework implements a corresponding three-level control hierarchy operating at different frequencies:
  • First-order level: Comprising the spinal cord module, this level operates at the highest frequency. Its primary function is to generate real-time control torques sent to the robotic arm, ensuring accurate and rapid responses.
  • Second-order level: Including the brainstem and thalamus module, this level operates at an intermediate frequency. It serves as a relay between the first- and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order loop and integrating and filtering the torques provided by the third-order layer to supply feedforward torque to the first-order loop.
  • Third-order level: Comprising the cerebellum and cerebral cortex module, this layer operates at the lowest frequency. Its role is to generate movement trajectories and provide gravity compensation torques based on motion patterns, maintaining movement stability and smoothness.
II-C Information pathway
As illustrated in Fig.1a, the human central nervous system employs ascending and descending pathways for sensory information transmission and motor command execution, respectively. The ascending pathways primarily convey peripheral sensory information to the brain for perception and integration, while the descending pathways transmit motor commands from the brain to the spinal cord and peripheral nerves to control movements. Following this principle, our framework incorporates two analogous information pathways:
  • Ascending pathway: Inspired by the dorsal column–medial lemniscal and spinocerebellar tracts[william_ascending_pathway], this pathway transmits proprioceptive and touch information, namely joint positions, velocities, and end-effector torque. It simulates the process by which muscle receptors relay signals to the CNS for processing and can be regarded as the input process of the framework.
  • Descending pathway: Inspired by the corticospinal and reticulospinal tracts[william_ascending_pathway], this pathway transmits joint torque commands and parameter adjustment signals. It simulates the process by which the CNS delivers motor commands to muscle effectors and can be regarded as the output process of the framework.
II-D A complete operational cycle
To clarify the interplay among modules, levels, and pathways, we illustrate in Fig. 1b a complete working cycle of the proposed framework, showing how sensory inputs are transformed into motor control outputs.
In each cycle, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$) and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$ are transmitted via first-order neurons to the spinal cord. Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem, which adaptively provides PD gains $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum, which outputs gravity compensation matrix $\bm{T}_{\text{cb}}$. These ascending transmissions complete the flow of proprioceptive information from sensors to the modules.
On the descending side, $\bm{T}_{\text{cb}}$ is conveyed via third-order neurons to the thalamus, where end-effector torque feedback $\bm{F}_{\text{ef}}$ is integrated to generate a weight vector $\bm{w}$. This vector filters $\bm{T}_{\text{cb}}$ to produce the effective gravity compensation torque $\bm{\tau}_{g}$, which is passed via second-order neurons to the spinal cord. There, $\bm{\tau}_{g}$ is combined with $\bm{\tau}_{\text{fb}}$ to form the final control torque $\bm{\tau}$, which is delivered through first-order neurons to the actuators. Thus, the descending pathway completes motor command transmission from the modules to the joints.
Because information reaches different modules through pathways of varying length, the framework is naturally organized into three control levels. The first-order level involves only spinal processing and yields the shortest latency and highest frequency. The second-order level adds brainstem and thalamic processing, resulting in longer pathways and lower frequency. The third-order level incorporates the cerebellum, producing the longest transmission path and the lowest frequency.
The pseudocode implementation of this framework is provided in Algorithm 1.
Algorithm 1 CBMC-V3
0:  joint proprioceptive sensor data $\bm{\dot{q}},\bm{q}$, planned trajectory data $\bm{\dot{q}}_{d},\bm{q}_{d}$, end-effector torque sensor data $\bm{F}_{ef}$
0:  joint control torque $\bm{\tau}$
1:  Main thread:
2:  initialize first-order, second-order and third-order level control cycle $T_{1},T_{2},T_{3}$
3:  initialize trajectory variables $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
4:  initialize end-effector torque variable $\bm{F}_{ef}$
5:  initialize control torque variable $\bm{\tau}$
6:  initialize intermediate variables $\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g},\bm{w},\bm{T}_{{\text{cb}}}$
7:  read trajectory length $M$
8:  for $t=1$ to $M$ do
9:   read data from sensors and update $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q},\bm{F}_{ef}$
10:   if $t$ Mod $T_{1}==0$ then
11:    run sub-thread 1
12:   end if
13:   if $t$ Mod $T_{2}==0$ then
14:    run sub-thread 2
15:    run sub-thread 3
16:   end if
17:   if $t$ Mod $T_{3}==0$ then
18:    run sub-thread 4
19:   end if
20:   send $\bm{\tau}$ to actuators
21:  end for
22:
23:  Sub-thread 1:
24:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q},\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g}$
25:  process through the spinal cord module
26:  update $\bm{\tau}$
27:
28:  Sub-thread 2:
29:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
30:  process through the brainstem module
31:  update $\bm{K}_{V},\bm{K}_{P}$
32:
33:  Sub-thread 3:
34:  read $\bm{T}_{{\text{cb}}},\bm{F}_{ef}$
35:  process through the thalamus module
36:  update $\bm{\tau}_{g},\bm{w}$
37:
38:  Sub-thread 4:
39:  read $\bm{q}$
40:  process through the cerebellum module
41:  update $\bm{T}_{{\text{cb}}}$
III Module network design
Building on the framework introduced in the previous section, this chapter focuses on the implementation of each module using SNNs. The cerebral cortex module, which serves only for trajectory storage and playback, is therefore excluded from the discussion. Since high-performance robotic arms are commonly designed with seven DoF, all implementations in this chapter target a 7-DoF arm. For manipulators with different DoF, the same design principles can be applied by appropriately scaling the neural network architecture of each module.
III-A Spiking neuron
SNN employs spiking neurons as their fundamental computational units. A variety of spiking neuron models have been proposed in the literature, such as the Hodgkin–Huxley model [hodgkin1952quantitative] and the Leaky Integrate-and-Fire (LIF) model along with its derivatives [burkitt2006review]. The LIF model is of particular interest here, since it not only captures the dynamic process of membrane potential accumulation and leakage, but also maintains relatively high computational efficiency [stein1965theoretical]. Therefore, we will focus on the LIF model.
The input current signal generated by a spike train $S(t)=\sum_{f}s(t-t_{f}),f=1,2,\cdots$ acting on a neuron can be described as
$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)exp\left(-s/\tau_{s}\right)ds}$ (7)
where $\tau_{s}$ denotes the synaptic time constant. The membrane potential $u$ of the neuron evolves according to the following dynamics:
$\tau_{m}\frac{du}{dt}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
where $\tau_{m}=RC$ is the membrane time constant determined by the membrane resistance $R$ and capacitance $C$. Here, $u_{\text{reset}}$ is the reset potential after firing, $i_{0}(t)$ represents the external current driving the neuron, $i_{j}(t)$ denotes the input current from the $j$-th synapse, and $w_{j}$ is the synaptic weight associated with the $j$-th synapse. When the membrane potential $u$ reaches a certain threshold $u_{\text{fire}}$, the neuron emits a spike and the potential is reset to $u_{\text{reset}}$. Therefore, the operational process of a spiking neuron can be interpreted as follows:
1.When the membrane potential $u(t)$ exceeds the firing threshold $u_{\text{fire}}$, the neuron emits a spike $s(t)$, expressed as
$s(t)=\Theta(u(t)-u_{\text{fire}})$ (9)
where
$\Theta(x)=\begin{cases}1,&x\geq 0\\ 0,&x<0\end{cases}$ (10)
is the Heaviside step function. Simultaneously, the membrane potential is reset to $u_{\text{reset}}$. When the membrane potential $u(t)$ remains below the threshold $u_{\text{fire}}$, no spike is generated.
2.At each time step, the differential equation (8) is updated, where the membrane potential decays by $\lvert u_{\text{reset}}-u(t)\rvert$ (representing the forgetting effect over time) and increases by $R\left(i_{0}(t)+\sum w_{j}i_{j}(t)\right)$ (representing the stimulation effect induced by external inputs).
3.The process then returns to Step 1.
In our framework, we adopt a backpropagation-based training method[8891809] for SNN. Since the activation function of the neuron in 10 is a discontinuous function, directly computing its derivative can lead to network instability. Consequently, various surrogate gradient methods have been proposed [8891809]. A common approach is to use the original step function $y_{i}=\Theta(a_{i})$ during the forward pass, while replacing its derivative with $\theta^{\prime}(a_{i})=\sigma^{\prime}(a_{i})$ instead of $\Theta^{\prime}(a_{i})$ during backpropagation, where $\sigma(a_{i})$ is referred to as the surrogate activation function. Typically, $\sigma(a_{i})$ is a smooth and continuous function that resembles the shape of $\Theta(a_{i})$ while enabling stable gradient computation.
As illustrated in Fig.1c, the framework employs three types of neurons across all modules, represented in gray, yellow, and green. The gray neurons correspond to input fibers, which serve as the data interface for each module via connections to the ascending or descending pathway. The yellow neurons represent LIF neurons, which constitute the core units of SNNs and are responsible for information storage, processing, and learning. The green neurons correspond to non-spiking LIF neurons, implemented by setting the membrane time constant $\tau_{m}$ of standard LIF neurons to infinity. Unlike spiking neurons, their output is the continuous membrane potential $u(t)$ rather than discrete spikes, thereby enabling spike-to-continuous signal conversion. These neurons serve as the output layer of each module, connecting to the descending pathway, and, in addition to the functional capabilities of the yellow LIF neurons, they also perform information decoding and output.
III-B Cerebellum module
Humans are capable of rapidly acquiring complex skills through imitation, and neural networks exhibit similar properties. Therefore, compared with unsupervised learning approaches, training a neural network to perform gravity compensation via imitation provides a faster and more efficient solution. Under the condition of a known robotic arm dynamics model, the gravity compensation torque represents a nonlinear mapping from joint positions to joint torques. In the machine learning domain, the task of fitting a curve to observed data under supervision is referred to as a regression task, for which extensive research and mature methods already exist[henkes2024spiking, robotics13090126].
Figure 2: (a) Structure of one network unit during training in the cerebellum module. (b) Structure of cerebellum network during prediction.
As shown in Fig.2a, to implement the cerebellum module, a four-layer recurrent SNN is designed as a network unit of the module, analogous to the “microcomplex” proposed in [abadia2019robot, abadia2021cerebellar]. The input variable $\bm{q}\in\mathbb{R}^{7}$ corresponds to the actual joint positions of the robotic arm, and the output $\bm{\tau}_{\text{cb}}\in\mathbb{R}^{7}$ represents the gravity compensation torques under a given load condition.
Variables from the ascending pathway first arrive at the mossy fiber (MF) layer, which is fully connected to the granule cell (GC) layer consisting of 300 LIF neurons. The GC layer, in turn, is fully connected to the Purkinje cell (PC) layer containing 200 LIF neurons. Each PC neuron receives input from both the current and previous time step of the GC layer, capturing temporal dynamics. The output of the PC layer is fully connected to the deep cerebellar nuclei (DCN) layer, which consists of 140 non-spiking LIF neurons. The decoding of the DCN layer is inspired by population coding observed in the motor cortex of animals [georgopoulos1986neuronal]. Studies have shown that during arm movement in primates, the activity of any single motor cortex neuron does not directly indicate the actual movement direction; however, a weighted sum of vectors corresponding to the preferred directions of a population of neurons yields a vector aligned with the movement direction. Analogously, the 140 DCN neurons are grouped in sets of 20, and their membrane potentials are averaged to obtain a 7-dimensional torque output $\bm{\tau}_{\text{cb}}$, i.e.,
$\tau_{\text{cb},i}=\sum_{j=20\cdot(i-1)}^{20\cdot i}u_{\text{DCN},j}$ (11)
where $\tau_{\text{cb},i},i=1,\cdots,7$ represents the torque for each joint and $u_{\text{DCN},j},j=1,\cdots,140$ denotes the membrane potential of each DCN neuron. The network employs rate-based encoding[gerstner2002spiking], with a time window of length $T$.
The training data $\bm{\tau}_{\text{ref}}$ are generated using the robotic arm’s dynamics model. Under the condition that the training data adequately cover the workspace, this regression-based approach is sufficient to achieve effective gravity compensation. The network is trained in a supervised manner, with the objective of minimizing the error between the torques output by the network and the reference torques computed from the dynamics model. Structurally, this is reflected in the DCN layer neurons receiving both excitatory inputs from the PC layer neurons and inhibitory inputs from the climbing fiber (CF) neurons. Mathematically, the network output torque $\bm{\tau}_{\text{cb}}$ is compared with the reference torque $\bm{\tau}_{\text{ref}}$ to obtain an error, which is quantified using the mean squared error (MSE Loss), i.e.,
$\text{MSE}=\frac{\sum_{i=1}^{7}(\tau_{\text{cb,i}}-\tau_{\text{ref,i}})^{2}}{7}$ (12)
The weights are then updated via backpropagation, thereby completing the training process; detailed procedures are provided in Algorithm 2.
Algorithm 2 Training process of one network unit in the cerebellum module
1:  initialize load $M$, trajectory length $L$, batch size bsbs, learning rate lrlr, momentum momo, epochs epep, random noise $\bm{n}$, network loss lossloss
2:  initialize reference torque $\bm{\tau}_{\text{ref}}[1:L]$, network output torque $\bm{\tau}_{\text{cb}}[1:bs]$
3:  initialize optimizer $optim(lr,mo)$
4:  initialize network
5:  load training trajectory $\bm{q}_{d}[1:L]$
6:  load robotic arm urdf file and initialize dynamic model according to $M$
7:  for $t=1$ to $L$ do
8:   read joint position $\bm{q}_{d}[t]$
9:   generate reference torque $\bm{\tau}_{\text{ref}}(\bm{q}_{d}[t])$ using dynamic model
10:   update $\bm{\tau}_{\text{ref}}[t]$
11:  end for
12:  for $i=1$ to epep do
13:   for $j=1$ to $L$ by bsbs do
14:    update random noise $\bm{n}$
15:    add noise to a batch of reference torque $\bm{\tau}_{\text{ref,n}}[j:j+bs]=\bm{\tau}_{\text{ref}}[j:j+bs]+\bm{n}$
16:    generate network output $\bm{\tau}_{\text{cb}}[j:j+bs]$
17:    calculate error between network output and reference torque $loss(\bm{\tau}_{\text{ref,n}}[j:j+bs],\bm{\tau}_{\text{cb}}[j:j+bs])$
18:    back propagate lossloss
19:    update network weights using optimoptim
20:   end for
21:  end for
During backpropagation, a surrogate gradient function is employed, given by
$\sigma(x)=\frac{1}{1+\exp^{-\nu x}}$ (13)
where $\nu$ is a scaling factor. In this study, horizontal circular and inclined circular trajectories are used for training, with their mathematical expressions defined as
horizontal circle: $\begin{cases}x=x_{0}+R_{c}\cos(\frac{2\pi t}{T_{c}})\\ y=y_{0}+R_{c}\sin(\frac{2\pi t}{T_{c}})\\ z=z_{0}\end{cases}$ (14)
inclined circle: $\begin{cases}x=x_{1}+R_{c}\cos(\frac{2\pi t}{T_{c}})\cos\theta_{0}\\ y=y_{1}+R_{c}\sin(\frac{2\pi t}{T_{c}})\\ z=z_{1}+R_{c}\cos(\frac{2\pi t}{T_{c}})\sin\theta_{0}\end{cases}$ (15)
The parameters of neurons in each layer, as well as other settings used during training, are summarized in Table I.
TABLE I: Training and neuron parameters of the cerebellum module
Training parameter Value Neuron parameter Value
Epochs 50 each trajectory $u_{\text{reset}}$ GC: 0.0
Learning rate 0.01 PC: 0.0
Momentum 0.5 DCN: 0.0
Noise $N(0,1)$ $u_{\text{fire}}$ GC: 0.1
Batch size 10 PC: 0.1
Optimizer SGD DCN: $\infty$
$\theta_{0}$ $\frac{\pi}{6}$ $\tau_{m}$ GC: 10.0
$R_{c}$ 0.14 PC: 10.0
$T_{c}$ 3 DCN: 5.0
$(x_{0},y_{0},z_{0})$ (0.54,0.0,0.45) $\nu$ 5.0
$(x_{1},y_{1},z_{1})$ (0.63,-0.11,0.3) $T$ 10
After training, the network can operate without reference torques, with CF neurons no longer required, relying solely on the learned synaptic weights to generate gravity compensation torques. To accommodate different loads, two network units are trained separately under light (no load) and heavy (3 kg load) modes using the supervised learning procedure described in Algorithm 2. These trained networks are then combined to form the complete cerebellum module structure illustrated in Fig. 2b, enabling adaptive gravity compensation across multiple load conditions.
In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$. This matrix corresponds to the learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy conditions. The outputs are subsequently integrated by the thalamus module to achieve gravity compensation for arbitrary loads within a specified range. During trajectory tracking, the prediction process of the cerebellum module is described in Algorithm 3, which can be implemented as Sub-thread 4 in Algorithm 1.
Algorithm 3 Prediction process of the cerebellum module
1:  load the network for light and heavy load conditions
2:  initialize gravity compensation matrix $\bm{T}_{\text{cb}}$
3:  read joint position $\bm{q}$
4:  feed $\bm{q}$ into the two trained network unit, and generate two output $\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}}$
5:  update $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$
III-C Thalamus module
As described in the previous sub-section, although we trained two neural network units under different end-effector load conditions to form the cerebellum module—thereby generating gravity compensation torques corresponding to two specific load cases—the cerebellum module alone remains insufficient for handling arbitrary load. To address this limitation, we developed the thalamus module. By utilizing torque sensor data at the robotic arm’s end-effector, the thalamus module network is able to “sense” the weight of the object, akin to human perception, and accordingly assign adaptive weights to the two gravity compensation torques produced by the cerebellum module’s motor patterns via reinforcement learning. This mechanism enables effective compensation for loads of varying mass and shape.
Figure 3: Structure of thalamus module.
The thalamus module is implemented as a three-layer fully connected SNN, as illustrated in Fig. 3. Its input, $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$, corresponds to the six-dimensional force/torque sensor data measured at the end-effector, while its output, $\bm{w}=(w_{1},w_{2})^{T}\in\mathbb{R}^{2}$, represents the weight coefficients assigned to the two gravity compensation torques provided by the cerebellum module.
The input layer is fully connected to a hidden layer consisting of five LIF neurons, which in turn connects to an output layer of two non-spiking LIF neurons. The membrane potentials of the output neurons are then passed through a Softmax function to produce the two weight coefficients. The loss function designed for this network comprises two terms, minimizing both the joint position error and the joint velocity error, i.e.,
$\text{Loss}=\sum_{i=1}^{7}100*(q_{d,i}-q_{i})+\sum_{i=1}^{7}10^{-4}\cdot(\dot{q}_{d,i}-\dot{q}_{i})$ (16)
Similar to the cerebellum module, the thalamus module adopts rate-based encoding with a time window length of $T$, and employs the sigmoid function as the surrogate gradient function.
TABLE II: Training and neuron parameters of the thalamus module
Parameter Value Parameter Value
Learning rate 0.01 $\nu$ 5.0
$T$ 10
Weight decay 0.001 $u_{\text{reset}}$ Hidden layer: 0.0
Optimizer SGD Output layer: 10.0
Momentum 0.5 Filter layer: 0.0
$\tau_{m}$ Hidden layer: 5.0 $u_{\text{fire}}$ Hidden layer: 0.01
Output layer: 5.0 Output layer: $\infty$
Filter layer: 5.0 Filter layer: $\infty$
After obtaining $\bm{w}$ from the output layer, it is then combined with data from the connecting fiber (linking the thalamus module and the cerebellum module) to generate a weighted output via matrix multiplication, yielding
$\begin{split}\delta\bm{u}_{\text{ft}}(t)&=(\delta u_{\text{ft,1}}(t),\delta u_{\text{ft,2}}(t),\cdots,\delta u_{\text{ft,7}}(t))^{T}\\ &=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\cdot\bm{w}=w_{1}\bm{\tau}_{\text{cb,L}}+w_{2}\bm{\tau}_{\text{cb,H}}\end{split}$ (17)
This weighted result is not directly used as the output; instead, it is injected as an increment to the membrane potentials of seven non-spiking LIF neurons that serve as a filter layer. Specifically, the membrane potential of each filtering-layer neuron at time $t$ is given by
$u_{\text{ft},i}(t)$ $=u_{\text{ft},i}(t-1)-\frac{u_{\text{ft},i}(t-1)-u_{\text{reset}}}{\tau_{m}}$ (18)
$+\frac{\delta u_{\text{ft},i}(t)}{\tau_{m}},i=1,\cdots,7$
Finally we take the membrane potential of each filter-layer neuron at time $t$ as the gravity compensation torque, i.e., $\bm{\tau}_{g}=\bm{u}_{\text{ft}}(t)=(u_{\text{ft},1}(t),u_{\text{ft},2}(t),\cdots,u_{\text{ft},7}(t))$, which is then transmitted to the spinal cord module for further processing. The parameters of the neurons in each layer of this network, as well as the training parameters, are summarized in Table II. During trajectory tracking tasks, the learning procedure of the thalamus module follows Algorithm 4, which can be directly implemented as Sub-thread 3 in Algorithm 1.
Algorithm 4 Training process of the thalamus module
1:  initialize gravity compensation torque $\bm{\tau}_{g}$, weight coefficients $\bm{w}$, network loss lossloss, optimizer $optim(lr,mo)$
2:  read $\bm{T}_{{\text{cb}}},\bm{F}_{\text{ef}}$
3:  feed $\bm{F}_{\text{ef}}$ into the network and generate weight coefficients $\bm{w}$
4:  calculate $\bm{T}_{\text{cb}}\cdot\bm{w}$ and feed the result into the filter layer neurons
5:  read the membrane potential of the filter layer $\bm{u}_{\text{ft}}$, update $\bm{\tau}_{g}$
6:  read joint position and velocity data $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
7:  calculate and update lossloss
8:  back propagate lossloss
9:  update network weights using optimoptim
III-D Brainstem module
As illustrated in Fig. 1c, the primary role of the brainstem module is to dynamically select appropriate feedback control coefficients during operation through learning, thereby compensating for the nonlinear dynamics of the robotic arm and the coupling effects among joints. The network is implemented as a three-layer fully connected SNN. Its input variable $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ is a concatenated vector consisting of the desired joint positions and velocities, as well as the actual joint positions and velocities. The output $\bm{K}=(k_{p,1},k_{p,2},\cdots,k_{p,7},k_{v,1},k_{v,2},\cdots,k_{v,7})\in\mathbb{R}^{14}$ corresponds to the proportional and derivative coefficients for each of the seven joints.
The input vector is fully connected to a hidden layer of 10 LIF neurons, which in turn is fully connected to an output layer of 14 non-spiking LIF neurons. The network is trained using reinforcement learning, with the loss function defined as
[…]
In this formulation, the first term minimizes the position error and the second term minimizes the velocity error, ensuring accurate trajectory tracking, while the third term penalizes torque usage to reduce joint energy consumption and prevent abrupt fluctuations in joint torque. Through this design, the network is able to learn feedback control coefficients that balance trajectory-tracking accuracy with reduced joint wear and energy cost.
Similar to the cerebellum module, the brainstem module adopts rate–based encoding with a time window length of $T$, and the approximate gradient function is chosen as the sigmoid function.
After decoding through the non-spiking LIF neurons in the output layer, we obtain a set of membrane potentials $\bm{u}_{\text{bs}}=(u_{\text{bs},1},\cdots,u_{\text{bs},14})\in\mathbb{R}^{14}$. By applying a scaling transformation, these membrane potentials can be mapped to the proportional and derivative coefficients, i.e.,
$\begin{split}\bm{K}_{P}&=diag(k_{p,1},k_{p,2},\cdots,k_{p,7})\\ &=a\cdot diag(u_{\text{bs,1}},u_{\text{bs,2}},\cdots,u_{\text{bs,7}})\end{split}$ (20)
$\begin{split}\bm{K}_{V}&=diag(k_{v,1},k_{v,2},\cdots,k_{v,7})\\ &=b\cdot diag(u_{\text{bs,8}},u_{\text{bs,9}},\cdots,u_{\text{bs,14}})\end{split}$ (21)
where $a$ and $b$ are two scaling parameters. The parameters of the neurons in each layer of the network, as well as the parameters used during training, are summarized in Table III.
TABLE III: Training and neuron parameters of the brainstem module
Parameter Value Parameter Value
Learning Rate 0.01 $u_{\text{reset}}$ Hidden layer: 0.0
Momentum 0.5 Output layer: 10.0
Optimizer SGD $u_{\text{fire}}$ Hidden layer: 0.005
$\nu$ 5.0 Output layer: $\infty$
(a,b) (20,2) $\tau_{m}$ Hidden layer: 5.0
$T$ 15 Output layer: 5.0
During trajectory tracking tasks, the learning process of the brainstem module is described in Algorithm 5, which can be implemented as Sub-thread 2 within Algorithm1.
Algorithm 5 Training process of the brainstem module
1:  initialize coefficients $\bm{K}_{V},\bm{K}_{P}$, network loss lossloss, optimizer $optim(lr,mo)$
2:  read joint position and velocity data $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
3:  feed $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]$ into the network
4:  read the potential of the output layer neurons $\bm{u}_{\text{bs}}$ and scale them through scaling parameters
5:  update $\bm{K}_{V},\bm{K}_{P}$
6:  read joint position and velocity data $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
7:  calculate and update lossloss
8:  back propagate lossloss
9:  update network weights using optimoptim
III-E Spinal cord module
The primary function of the spinal cord module is to integrate proprioceptive information with the outputs of the aforementioned higher-level motor control modules, thereby generating real-time motion control commands. Its structural design is illustrated in Fig. 1c. For high-performance robotic manipulators, torque control typically requires the control algorithm to achieve frequencies of 1000 Hz, 1200 Hz, or even higher. Consequently, the spinal cord module must exhibit very high computational efficiency.
In our design, the input variables $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ consist of the planned positions and velocities as well as the actual positions and velocities of all joints, while the output $\bm{\tau}\in\mathbb{R}^{7}$ represents the control torque for each joint. After entering the spinal cord module, the planned trajectory and proprioceptive feedback are first encoded into discrete spike trains, with each joint being processed independently. Specifically, the planned position and velocity signals $q_{d,i},\dot{q}_{d,i},i=1,\cdots,7$ serve to excite the input fibers, whereas the actual proprioceptive feedback $q_{i},\dot{q}_{i},i=1,\cdots,7$ inhibits the input fibers. Mathematically, this process is equivalent to computing their difference.
The outputs of the input fibers are subsequently passed into the Poisson encoding stage, given by
$s_{q,i,j}=\begin{cases}0,&-u_{j}<\frac{q_{d,i}-q_{i}}{\Delta q_{i}}<u_{j},j=1,\cdots,n\\ 1,&\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=1,\cdots,\frac{n}{2}\\ -1,&-\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=\frac{n}{2},\cdots,n\end{cases}$ (22)
$s_{\dot{q},i,j}=\begin{cases}0,&-u_{j}<\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}<u_{j},j=1,\cdots,n\\ 1,&\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=1,\cdots,\frac{n}{2}\\ -1,&-\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=\frac{n}{2},\cdots,n\end{cases}$ (23)
Here, $u_{j}\sim U(0,1),j=1,\cdots,n$ denotes $n$ independent uniformly distributed random variables, ${s_{q,i,j}},i=1,\cdots,7,j=1,\cdots,n$ represents the spike trains encoding the position information of each joint at the current time step, and ${s_{\dot{q},i,j}},i=1,\cdots,7,j=1,\cdots,n$ denotes the spike trains encoding the velocity information. The terms $\Delta q_{i}$ and $\Delta\dot{q}_{i},i=1,\cdots,7$ correspond to the maximum allowable position and velocity errors for each joint, respectively.
From an intuitive perspective, each input fiber is connected to a group of $n$ submodules, where the first half $\frac{n}{2}$ are responsible for processing positive inputs and the second half $\frac{n}{2}$ handle negative inputs. This structure corresponds to the agonist–antagonist pairing characteristic of human muscles. Each submodule generates a uniformly distributed random number in the range $[0,1]$. When the output of the input fiber is positive, the first $\frac{n}{2}$ neurons are encoded sequentially: if the absolute value of the input exceeds the random number, the encoding result is set to $1$; otherwise, it is 0. Conversely, when the output of the input fiber is negative, the second $\frac{n}{2}$ neurons are encoded sequentially: if the absolute value of the input exceeds the random number, the encoding result is set to 1-1; otherwise, it is 0. This encoding scheme can be implemented in hardware using simple logic gate circuits, thereby achieving very high computational efficiency.
The encoding precision is influenced by three parameters: $\Delta q_{i}$, $\Delta\dot{q}_{i}$, $i=1,\cdots,7$, and $n$. Smaller values of $\Delta q_{i}$ and $\Delta\dot{q}_{i}$ lead to higher encoding precision, but reduce the range of representable errors; a larger $n$ increases encoding precision but reduces computational efficiency.
Following this encoding process, two sets of discrete spike trains are generated—one representing position information and the other representing velocity information. These spike trains are then transmitted to stimulate the activity of non-spiking LIF neurons in the output layer, with stimulation strength modulated by the feedback control gains $k_{p,i}$ and $k_{v,i}$ provided by the brainstem module, as illustrated by the blue and red dashed lines, respectively. Accordingly, at each time step, the membrane potential increments of the output non-spiking LIF neurons can be expressed as
$\delta u_{\text{sp,i}}(t)$ $=\tau_{g,i}+k_{p,i}\cdot\sum_{j=1}^{100}s_{q,i,j}$ (24)
$+k_{v,i}\cdot\sum_{j=1}^{100}s_{\dot{q},i,j},i=1,\cdots,7$
Here, $\bm{\tau}_{g}=(\tau_{g,1},\tau_{g,2},\cdots,\tau_{g,7})$ denotes the gravity compensation torque generated by the thalamus module. Consequently, the membrane potentials at each time step can be written as
$u_{\text{sp},i}(t)$ $=u_{\text{sp},i}(t-1)-\frac{u_{\text{sp},i}(t-1)-u_{\text{reset}}}{\tau_{m}}$ (25)
$+\frac{\delta u_{\text{sp},i}(t)}{\tau_{m}},i=1,\cdots,7$
Finally, similar to the thalamus module, we take $\bm{\tau}=\bm{u}_{\text{sp}}=(u_{\text{sp},1}(t),u_{\text{sp},2}(t),\cdots,u_{\text{sp},7}(t))$ as the final control torque, which will be sent in real time to the actuators. The neuronal parameters of the spinal cord module are summarized in Table IV.
TABLE IV: Neuron parameters of the spinal cord module
Parameter Value Parameter Value
$u_{\text{reset}}$ 0.0 $\Delta q_{i},i=1,\cdots,7$ 0.5
$u_{\text{fire}}$ $\infty$ $\Delta\dot{q}_{i},i=1,\cdots,7$ 0.5
$\tau_{m}$ 2.0 n 100
Through this mechanism, the output torque at each time step incorporates both the current information and the historical information over a preceding time window, thereby making it more suitable than ANN for time-continuous tasks such as robotic control. During trajectory tracking tasks, the operational process of the spinal cord module is illustrated in Algorithm 6, which can be implemented as Sub-thread 1 within Algorithm1.
Algorithm 6 Prediction process of the spinal cord module
1:  initialize control torque $\bm{\tau}$
2:  read joint position and velocity data $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
3:  read proportional and derivative coefficients $\bm{K}_{V},\bm{K}_{P}$ and gravity compensation torque $\bm{\tau}_{g}$
4:  poisson encode $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$ to generate two sets of spike trains $s_{q,i,j}],[s_{\dot{q},i,j}]$
5:  modulate the spike trains with $\bm{K}_{V},\bm{K}_{P}$
6:  feed $\bm{\tau}_{g}$ and the modulated spike trains into the non-spiking LIF neurons
7:  read potential and update $\bm{\tau}$
IV Simulation analysis
To evaluate the rationality and effectiveness of the proposed framework, we conduct three simulation experiments that demonstrate contributions (1)–(3). In section IV-A and IV-B, we validate the design of the modules and in section IV-C, the design of the framework.
The algorithm is implemented in Python (version 3.12.4), and all neural networks are constructed using PyTorch [paszke2019pytorch] (version 2.4.0). The spiking neurons involved in the algorithm are implemented with the SNN learning framework SpikingJelly [doi:10.1126/sciadv.adi1480] (version 0.0.0.0.14). The simulation environment employed is PyBullet [coumans2021] (version 3.2.6), an open-source robotics simulation tool whose underlying physics engine is based on the open-source Bullet Physics SDK. The controlled robotic arm consists of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector. The simulation time step is set to 1 ms, and the control frequency is configured to 1 kHz.
In all simulation experiments presented in this section, trajectory tracking accuracy is evaluated using the root-mean-square error (RMSE) of the joint positions. The RMSE at each time step is defined as
$\text{RMSE}(t)=\sqrt{\sum_{i=1}^{7}(q_{d,i}(t)-q_{i}(t))^{2}}$ (26)
while the average RMSE over an entire trajectory is defined as
$\text{RMSE}=\frac{1}{N}\sum_{t=1}^{N}RMSE(t)$ (27)
Three types of trajectories are employed in the simulations: horizontal circle, inclined circle, and spatial figure-eight. The first two trajectories have already been defined in section III-B, while the third trajectory is defined as
spatial figure-eight: $\begin{cases}x=x+0.5R\sin(\frac{4\pi t}{T})\\ y=y+R\cos(\frac{2\pi t}{T})\\ z=z+z\sin(\frac{2\pi t}{T})\end{cases}$ (28)
where $x=0.61,y=0,z=0.3,R=0.14,T=3,z=0.08$.
Figure 4: Average feedback control coefficients of each joint across 3 trajectory tracking tasks.
IV-A Spinal cord and brainstem module validation
Physiological studies have demonstrated that joint stiffness and damping in the human arm exhibit an inverse correlation [bennett1992time]. Specifically, stiffness decreases when a joint begins to move and increases when at rest, while damping decreases as the joint approaches a target and increases at the onset of motion. In the proposed framework, the feedback control coefficients $\bm{K}_{V}$ and $\bm{K}_{P}$ generated by the brainstem module correspond precisely to damping and stiffness in feedback control. Fig. 4 illustrates the variations of average $k_{p}$ (blue solid lines) and $k_{v}$ (red dashed lines) for each joint across three trajectories. It can be observed that the two parameters generally exhibit opposite trends, as exemplified by joints 3–6 during 3–5 s and joints 4 and 6 during 1–2 s. This finding indicates that the brainstem module, through reinforcement learning, adjusts joint stiffness and damping in the spinal module in a manner that exhibits biomimetic characteristics, consistent with our design objectives. We also note that this feature is less pronounced in joints closer to the robot’s base, such as joints 1 and 2, likely due to stronger coupling effects from adjacent joints.
Figure 5: (a) Average tracking error using single mode and two modes across three trajectories. (b) Average weight generated by the thalamus module across three trajectories.
IV-B Thalamus and cerebellum module validation
Fig. 5a shows that single gravity compensation modes from the cerebellum module are only effective within narrow load ranges (e.g., the light mode around 1 kg and the heavy mode around 3.5 kg), with errors otherwise increasing nearly linearly with load variation. In contrast, the thalamus module adaptively synthesizes both modes, maintaining stable joint error across the entire 1–3.5 kg range, thereby validating its design. As shown in Fig. 5b, reinforcement learning enables the thalamus module to adjust mode weights according to load mass, assigning a smaller weight to the light mode and a greater weight to the heavy mode as the load increases. This adaptive mechanism mirrors human motor control, where movements are achieved through linear combinations of motor primitives[sanger2000human], thus highlighting the biomimetic nature of the thalamus and cerebellum module.
Figure 6: Average joint error across the three trajectories under varying load conditions during operation. (a) 0-1.5s. (b) 12-14s. (c) 24-25.5s. (d) 0-1.5s without the brainstem module. (e) 0-1.5s without the thalamus module. (f) 0-1.5s without the cerebellum module.
IV-C Framework validation
Fig. 6 illustrates the average trajectory tracking error across three trajectories under different load conditions. As shown in Fig. 6a, when a 1.5kg load is applied at 0s, the algorithm converges to a stable state within approximately 1.5s; after 12s of operation, adding an additional 2kg load again leads to convergence within about 2s; and upon removing the added load at 24s, convergence is also achieved within 2s. In contrast, the cerebellum-inspired network proposed in [abadia2019robot, abadia2021cerebellar] requires nearly 100s to adapt to a 500g change in load, highlighting the superior dynamic performance and convergence speed of our framework. Furthermore, we observe that, compared with cold-start initialization, subsequent load changes during ongoing operation induce smaller error fluctuations, as reflected by reduced peak magnitude at 12s and 24s, which can be attributed to the membrane potential dynamics of spiking neurons that preserve historical information and facilitate smoother adaptation.
Fig. 6b, 6c and 6d further show that the convergence process consistently exhibits three distinct error peaks following each load change. These peaks directly correspond to the hierarchical loop structure of the framework: the spinal cord module (first-order level) updates at the control frequency and therefore corresponds to every point on the curve; the brainstem and thalamus modules (second-order level) yield the earlier two peaks which reflect the learning convergence process; and the cerebellum module (third-order level) produces the latest peak owing to its longest processing cycle. The temporal separation of these peaks validates the layered organization of the control framework.
To verify this correspondence, ablation studies were conducted by selectively disabling individual modules, as shown in Fig. 6e, 6f and 6g. Removing the brainstem’s learning capability by fixing feedback control coefficients reduces the number of peaks to two and delays the first peak, indicating diminished adaptability to load change. Similarly, disabling the thalamus module by fixing mode-combination weights eliminates its characteristic peak, leaving only brainstem and cerebellar responses. Finally, replacing the cerebellum’s adaptive torque generation with fixed compensation suppresses the cerebellar peak and weakens the thalamic one, since the thalamus relies on cerebellar outputs to optimize weight assignment. These results confirm the functional contributions of each module and their hierarchical cooperation.
In summary, the proposed framework exhibits robust adaptability and rapid convergence to load variations, with distinct loop-specific responses emerging in a biologically consistent sequence. This layered activation mirrors the progression of human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [forgaard2015voluntary], thereby reinforcing the plausibility and effectiveness of the neuromorphic design.
V Experiment results
This section will present real-world experiments on a robotic arm platform to demonstrate contribution (4). All spiking neurons, except those in the spinal cord module, are implemented in Python using an open-source SNN learning framework, while the spinal cord module is realized in C++ through direct modeling of membrane potential dynamics. The deep learning framework, SNN framework, and computing hardware are identical to that used in the simulation. The robotic arm platform is a Flexiv Rizon 4s equipped with a Flexiv GRAV gripper, as shown in Fig. 7.
Figure 7: The experiment setting.
The experiments are conducted under nine load conditions, including two material types (sand-filled and water-filled containers) and five mass levels ranging from 1.5 kg (gripper with empty container) to 3.5 kg (container with 2 kg load). The upper limit is chosen below the 4 kg load capacity of the robotic platform to maintain a safety margin.
Eight end-effector trajectories (T1–T8, defined in the Appendix) are tested, all with vertically downward orientation and origin at the base flange center. Among them, T1–T5 represent slower, lower-dynamic trajectories, while T6–T8 correspond to faster, higher-dynamic ones. For statistical validity, each trajectory–load–controller combination is repeated 10 times. In all experiments presented in this section, the trajectory tracking effect is evaluated using the RMSE of the end-effector position and orientation. At each time step, the position and orientation RMSE are defined as
$RMSE_{p}(t)=$ (29)
$\sqrt{(x_{d}(t)-x(t))^{2}+(y_{d}(t)-y(t))^{2}+(z_{d}(t)-z(t))^{2}}$
$RMSE_{o}(t)=$ (30)
$\sqrt{(\theta_{d}(t)-\theta(t))^{2}+(\psi_{d}(t)-\psi(t))^{2}+(\phi_{d}(t)-\phi(t))^{2}}$
where $x_{d},y_{d},z_{d},\theta_{d},\psi_{d},\phi_{d}$ denote the desired end-effector position in Cartesian space and the desired pitch, yaw, and roll angles, respectively, and $x,y,z,\theta,\psi,\phi$ denote the corresponding actual values. The average position and orientation RMSE of a trajectory are defined as
$RMSE_{p}=\frac{1}{N}\sum_{t=1}^{N}RMSE_{p}(t)$ (31)
$RMSE_{o}=\frac{1}{N}\sum_{t=1}^{N}RMSE_{o}(t)$ (32)
where $N$ represents the trajectory length.
Figure 8: Trajectory tracking result under different trajectory and load conditions. (a) Position RMSE under sand load. (b) Orientation RMSE under sand load. (c) Position RMSE under water load. (d) Orientation RMSE under water load.
Fig. 8a and 8b show the trajectory tracking errors of the robotic arm when the end-effector container is filled with sand, under various trajectory and load conditions. Red and blue bars represent the proposed method and the factory-installed industrial-grade Flexiv Cartesian Position Control, respectively. Cross-hatched and slanted-hatched bars indicate the error ranges of trajectories T1–T5 (lower dynamic group) and T6–T8 (higher dynamic group) for a given load, while error bars denote the extrema of the errors. Four observations can be drawn:
1. Higher tracking accuracy: For any load mass, the proposed method yields smaller mean position and orientation errors, as indicated by the consistently lower red bars compared with the blue bars. Quantitatively, the Flexiv method achieves average errors of $1.789\times 10^{-2}$ m (position) and $1.685\times 10^{-2}$ rad (orientation) across all trajectories and load conditions, whereas the proposed method achieves $1.447\times 10^{-2}$ m and $1.473\times 10^{-2}$ rad, corresponding to reductions of 19.1% and 12.6%, respectively. These results demonstrate that the proposed method achieves higher control accuracy, thereby exhibiting superiority in the “Precision” indicator of agility.
2. Adaptation to load variation: As load mass increases, position errors of the proposed method gradually increase while orientation errors decrease, and both remain generally lower than those of the Flexiv method, as shown by the upward trend of red bars in Fig. 8a and the downward trend in Fig. 8b. This observation is consistent with the simulation results, confirming that the proposed framework can adapt to arbitrary load masses within a practical range, validating the design of the cerebellar and thalamic modules.
3. Consistency across trajectories: Within each trajectory group of similar dynamics, the proposed method exhibits greater consistency across different trajectory shapes, as the red bars for both hatching styles are shorter than their blue counterparts. Specifically, the variation ranges of Flexiv errors in position and orientation are $0.447\times 10^{-2}$ m and $0.592\times 10^{-2}$ rad, while those of the proposed method are $0.152\times 10^{-2}$ m and $0.363\times 10^{-2}$ rad, corresponding to reductions of 66.0% and 38.7%. This aligns with simulation findings, indicating that the proposed method can rapidly adapt online to arbitrary motion patterns, validating the brainstem and spinal cord modules and demonstrating an advantage in the agility indicator of “Transitions between diverse motion patterns”.
4. Performance under higher dynamics: For the higher-dynamic group, the performance improvement of the proposed method is more pronounced, as the gap between red and blue slanted bars is larger than that between the cross-hatched bars. Quantitatively, in the higher-dynamic group, the proposed method reduces position and orientation errors by 24.7% and 17.0% relative to the Flexiv method, exceeding the average improvements reported in conclusion 1. This outcome is expected, since the Flexiv method, being model-based, necessarily involves linear simplifications to ensure computational efficiency. While such simplifications suffice at low speeds where nonlinear effects (e.g., Coriolis and centrifugal forces) are weak, they deteriorate at higher speeds. In contrast, the proposed method adapts online to nonlinear dynamics, leading to smaller error increases. These results suggest that the proposed framework is more suitable for fast motions, demonstrating superiority in the agility indicator of “Speed” .
Fig. 8c and 8d present results when the container is filled with water. Compared to sand, water introduces stronger dynamic variations, thereby exerting greater influence on control performance. The proposed method yields mean errors of $1.442\times 10^{-2}$ m (position) and $1.519\times 10^{-2}$ rad (orientation), representing increases of 0.9% and 2.0% relative to the sand condition; the Flexiv method yields $1.779\times 10^{-2}$ m and $1.681\times 10^{-2}$ rad, representing decreases of 1.2% and 0.7%, respectively. Both methods show comparable trends across different load materials, indicating the ability to adapt to varying material properties. Moreover, the proposed method consistently demonstrates similar control performance under sand and water conditions, and analogous conclusions to the four points above can be drawn, which are omitted here for brevity.
In summary, the experiment results demonstrate that the proposed framework consistently outperforms the industrial-grade control algorithm across the three key indicators of agility—precision, adaptability to diverse motion patterns, and speed as mentioned in section I—thereby highlighting its effectiveness and the enhanced agility capacity provided by the brain-inspired design.
VI Conclusion and future works
In this paper, we propose a CNS-inspired control framework based on SNN, enabling agile robotic arm control in complex environments. The framework introduces a hierarchical architecture with five modules, is fully implemented using SNNs, and has been validated through both simulations and real-world robotic experiments, outperforming an industrial-grade baseline. Future work will focus on improving reinforcement learning for higher reproducibility, enhancing the cerebellar module for better gravity compensation, and further increasing control accuracy.
[Experiment trajectory setting]
$\text{T1:}\begin{cases}x=x_{0}+R_{0}\cos(\frac{2\pi t}{T_{0}})\\ y=y_{0}+R_{0}\sin(\frac{2\pi t}{T_{0}})\\ z=z_{0}\end{cases}\text{T2:}\begin{cases}x=x_{1}\\ y=y_{1}+R_{1}\sin(\frac{2\pi t}{T_{1}})\\ z=z_{1}+R_{1}\cos(\frac{2\pi t}{T_{1}})\end{cases}$
$\text{T3:}\begin{cases}x=x_{2}+R_{2}\cos(\frac{2\pi t}{T_{2}})\cos\theta_{1}\\ y=y_{2}+R_{2}\sin(\frac{2\pi t}{T_{2}})\\ z=z_{2}+R_{2}\cos(\frac{2\pi t}{T_{2}})\sin\theta_{1}\end{cases}$
$\text{T4:}\begin{cases}x=x_{3}+0.5R_{3}\sin(\frac{4\pi t}{T_{3}})\\ y=y_{3}+R_{3}\cos(\frac{2\pi t}{T_{3}})\\ z=z_{3}+z\sin(\frac{2\pi t}{T_{3}})\end{cases}$
$\text{T5:}\begin{cases}x=x_{3}+0.5R_{3}\sin(\frac{4\pi t}{T_{4}})\\ y=y_{3}+R_{3}\cos(\frac{2\pi t}{T_{4}})\\ z=z_{3}+z\sin(\frac{2\pi t}{T_{4}})\end{cases}$
$\text{T6:}\begin{cases}x=x_{2}+R_{2}\cos(\frac{2\pi t}{T_{5}})\cos\theta_{1}\\ y=y_{2}+R_{2}\sin(\frac{2\pi t}{T_{5}})\\ z=z_{2}+R_{2}\cos(\frac{2\pi t}{T_{5}})\sin\theta_{1}\end{cases}$
$\text{T7:}\begin{cases}x=x_{4}+R_{4}\cos(\frac{2\pi t}{T_{6}})\cos\theta_{2}\\ y=y_{4}+R_{4}\sin(\frac{2\pi t}{T_{6}})\\ z=z_{4}+R_{4}\cos(\frac{2\pi t}{T_{6}})\sin\theta_{2}\end{cases}$
$\text{T8:}\begin{cases}x=x_{5}+R_{5}\cos(\frac{2\pi t}{T_{7}})\cos\theta_{2}\cos\phi_{1}-R_{4}\sin(\frac{2\pi t}{T_{7}})\sin\phi_{1}\\ y=y_{5}+R_{5}\cos(\frac{2\pi t}{T_{7}})\cos\theta_{2}\sin\phi_{1}+R_{4}\sin(\frac{2\pi t}{T_{7}})\cos\phi_{1}\\ z=z_{5}+R_{5}\cos(\frac{2\pi t}{T_{7}})\sin\theta_{2}\end{cases}.$
TABLE V: Trajectory parameters
Parameter Value Parameter Value Parameter Value
$(x_{0},y_{0},z_{0})$
(0.54,0.0,
0.45) m
$R_{0}$ 0.14 m $T_{0}$ 4 s
$(x_{1},y_{1},z_{1})$
(0.6,0.35,
0.0) m
$R_{1}$ 0.2 m $T_{1}$ 4 s
$(x_{2},y_{2},z_{2})$
(0.63,-0.11,
0.3) m
$R_{2}$ 0.14 m $T_{2}$ 4 s
$(x_{3},y_{3},z_{3})$
(0.0,0.61,
0.3) m
$R_{3}$ 0.14 m $T_{3}$ 4s
$(x_{4},y_{4},z_{4})$
(0.53,-0.11,
0.4) m
$R_{4}$ 0.2 m $T_{4}$ 3s
$(x_{5},y_{5},z_{5})$
(0.53,-0.11,
0.4) m
$R_{5}$ 0.2 m $T_{5}$ 2s
$\theta_{1}$ $\frac{\pi}{6}$ rad $\phi_{1}$ $\frac{\pi}{3}$ rad $T_{6}$ 3s
$\theta_{2}$ $\frac{\pi}{6}$ rad $T_{7}$ 3s
[Uncaptioned image] Yanbo Pang received the B.Eng. from Sun Yat-sen University in 2022 and the M.Sc. from Tsinghua University in 2025. He is currently a Ph.D student at the Technical University of Munich (TUM). His research interests include neuromorphic computing, neurorobotics and robotic arm control.
[…]

### v2

@@
 Abstract
-As robotic arm applications extend beyond industrial settings into healthcare, service, and daily life, existing control algorithms struggle to achieve the agile manipulation required for complex environments with dynamic trajectories, unpredictable interactions, and diverse objects.
+As robotic arm applications extend beyond industrial settings into service-oriented sectors such as catering, household and retail, existing control algorithms struggle to achieve the agile manipulation required for complex environments with dynamic trajectories, unpredictable interactions, and diverse objects.
 This paper presents a biomimetic control framework based on Spiking Neural Networks (SNN), inspired by the human Central Nervous System (CNS), to achieve agile control in such environments.
-The proposed framework features five control modules (cerebral cortex, cerebellum, thalamus, brainstem, spinal cord), three hierarchical control levels (first-order, second-order and third-order), and two information pathways (ascending and descending).
+The proposed framework features five control modules (cerebral cortex, cerebellum, thalamus, brainstem, and spinal cord), three hierarchical control levels (first-order, second-order, and third-order), and two information pathways (ascending and descending).
 Each module is fully implemented using SNN.
@@
 The brainstem module employs a network of LIF and non-spiking LIF neurons to dynamically adjust spinal cord parameters via reinforcement learning.
-The thalamus module similarly adjusts the cerebellum’s torque outputs.
-The cerebellum module uses a recurrent SNN to learn the robotic arm’s dynamics through regression, providing feedforward gravity compensation torques.
+The thalamus module similarly employs a network of LIF and non-spiking LIF neurons to adjusts the cerebellum’s torque outputs via reinforcement learning.
+The cerebellum module, which provides feedforward gravity compensation torques, uses a recurrent SNN to learn the robotic arm’s dynamics through regression.
 The framework is validated both in simulation and on real-world robotic arm platform under various loads and trajectories.
@@
 I Introduction
-With the advancement of actuator and material technologies, robotic arms have gradually increased in degrees of freedom while decreasing in cost, enabling them to perform more delicate operations.
-Consequently, their applications have expanded in recent years to fields such as healthcare, services, and entertainment.
-As robotic arms become increasingly integrated into our daily lives, the complexity of their operating environments and the agility required for diverse, fine-grained tasks pose new challenges to control strategies.
-For instance, a 7-DOF robotic arm can readily cut a circular contour on a workpiece or transport heavy objects, yet it still struggles to perform a task as subtle as peeling an apple—a task easily accomplished by the human arm.
-Compared with biological systems, even the most advanced robotic arms today fall short of the agility of the human upper limb [mason2018], underscoring the existence of advanced motor control mechanisms in humans that remain underutilized in robotic systems.
+In recent years, robotic arms are increasingly applied in service-oriented sectors such as catering, household and retail.
+Compared with traditional industrial settings, these scenarios feature more complex environments and diverse tasks while requiring better agility, thereby posing new challenges to control strategies.
+For instance, a 7-DOF robotic arm can readily cut a circular contour on a workpiece or transport heavy replacement parts, yet it still struggles to perform a task as subtle as peeling an apple, which can be easily accomplished by the human arm.
+Compared with biological systems, even the most advanced robotic arms today fall short of the agility of human upper limb [mason2018], underscoring the existence of advanced motor control mechanisms in humans that remain underutilized in robotic systems.
 In well-modeled environments, control algorithms combining Whole-Body Control (WBC) [Abe2007, liu2011interactive, weighttans2011, Lee2012Intermediate, Escande2014, Kim2019, han2021recursive] and Model Predictive Control (MPC) [carron2019data, pankert2020perceptive, wang2023trajectory] can effectively counteract disturbances and achieve strong control performance.
@@
 However, they lack the adaptability inherent to human motor learning.
-In complex scenarios, variations in the size, mass, or dynamics of manipulated objects may render the kinematic and dynamic models inaccurate, limiting the effectiveness of purely model-based strategies.
+In complex scenarios, variations in the size, mass, or materials of manipulated objects may render the kinematic and dynamic models inaccurate, limiting the effectiveness of these model-based strategies.
 With the rapid progress of neural networks, their powerful approximation and learning capabilities have motivated researchers to employ them for adaptive model correction in response to environmental changes.
 For example, [chaoui2009ann] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [he2017adaptive] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [wang2018adaptive] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [salloom2020adaptive] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [liu2019adaptive] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [pham2020adaptive] applied radial basis function networks with dynamic surface control for robust dual-arm coordination.
-While these approaches achieved progress, they rely on artificial neural networks (ANN), which inherently disregard the temporal-spatial characteristics of motion, focusing only on instantaneous control.
+Despite these progression, they rely on artificial neural networks (ANN), which inherently disregard the temporal-spatial characteristics of motion, focusing only on instantaneous control.
 Furthermore, they remain constrained by robotic dynamics models, with network structures and training methods often tailored to specific problems and tightly coupled to complex mathematical formulations, thereby limiting generalizability.
@@
 By emulating biological neurons, SNNs process and store information through spike trains and membrane potential dynamics, offering advantages in computational speed, energy efficiency, information dimension, and biological plausibility, making them well-suited for robotic control [bing2018survey].
-For instance, [carrillo2008real] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn manipulator dynamics for torque control; [abadia2019robot] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [abadia2021cerebellar] further established resilience to control delays, highlighting potential in teleoperation and cloud robotics and refines this framework to enable robotic arms to dynamicallyy adjust compliance.
-However, these methods suffer from high structural complexity, computational cost, hyperparameter sensitivity, and limited interpretability, as they focus narrowly on cerebellar functionality while neglecting the integrative role of the broader central nervous system (CNS), which coordinates control through interactions among several neural areas[Halassa_2022].
+For instance, [carrillo2008real] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn manipulator dynamics for torque control; [abadia2019robot] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [abadia2021cerebellar] further established resilience to control delays, highlighting potential in teleoperation and cloud robotics; [doi:10.1126/scirobotics.adp2356] refines this framework to enable robotic arms to dynamically adjust compliance.
+However, these methods suffer from structural complexity, computational cost, hyperparameter sensitivity, and limited interpretability, as they focus narrowly on cerebellar functionality while neglecting the integrative role of the broader central nervous system (CNS), which coordinates control through interactions among several neural areas[Halassa_2022].
 Inspired by this systemic perspective of the CNS, [li2023cbmc] proposed the CBMC framework, incorporating modules such as the cortex, cerebellum, and spinal cord.
@@
 The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137].
-Motivated by this, the brainstem module in our framework dynamically adjusts the PD gains for each joint based on trajectory information,i.e.,
+Motivated by this, the brainstem module in our framework dynamically adjusts the PD gains for each joint based on trajectory information, providing the spinal cord module with time-varying proportional and derivative coefficients, i.e.,
 $\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{v,1},k_{v,2},\cdots,k_{v,n})$ (2)
 $\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{p,1},k_{p,2},\cdots,k_{p,n})$ (3)
-providing the spinal cord module with time-varying proportional and derivative coefficients, i.e., $k_{p,i}$ and $k_{v,i}$ for $i=1,\dots,n$.
 By integrating spinal feedback control with brainstem-driven adaptive gain modulation, the framework circumvents the complexity of manipulator dynamics while compensating for structured uncertainties through learning.
@@
 $\bm{\tau}=\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})+\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{1},\cdots,\tau_{n})^{T}$ (6)
-where $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ is produced by the spinal cord and brainstem module, $\bm{\tau}g(\bm{q},\bm{F}_{\text{ef}})$ is produced by the cerebellum and thalamus module, and $(\tau_{1},\dots,\tau_{n})^{T}$ represents the joint torques delivered to the actuators.
+where $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ is produced by the spinal cord and brainstem module, $\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})$ is produced by the cerebellum and thalamus module, and $(\tau_{1},\dots,\tau_{n})^{T}$ represents the joint torques delivered to the actuators.
 In conclusion, the function of each module can be summarized as follows:
@@
   • Second-order level: Including the brainstem and thalamus module, this level operates at an intermediate frequency.
-It serves as a relay between the first- and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order loop and integrating and filtering the torques provided by the third-order layer to supply feedforward torque to the first-order loop.
+It serves as a relay between the first-order and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order level and integrating and filtering the torques provided by the third-order level to supply feedforward torque to the first-order level.
   • Third-order level: Comprising the cerebellum and cerebral cortex module, this layer operates at the lowest frequency.
@@
 1b a complete working cycle of the proposed framework, showing how sensory inputs are transformed into motor control outputs.
-In each cycle, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$) and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$ are transmitted via first-order neurons to the spinal cord.
+In each cycle, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$ and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$) are transmitted via first-order neurons to the spinal cord.
 Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem, which adaptively provides PD gains $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum, which outputs gravity compensation matrix $\bm{T}_{\text{cb}}$.
@@
 The input current signal generated by a spike train $S(t)=\sum_{f}s(t-t_{f}),f=1,2,\cdots$ acting on a neuron can be described as
-$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)exp\left(-s/\tau_{s}\right)ds}$ (7)
+$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)\exp\left(-s/\tau_{s}\right)\text{d}s}$ (7)
 where $\tau_{s}$ denotes the synaptic time constant. The membrane potential $u$ of the neuron evolves according to the following dynamics:
-$\tau_{m}\frac{du}{dt}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
+$\tau_{m}\frac{\text{d}u}{\text{d}t}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
 where $\tau_{m}=RC$ is the membrane time constant determined by the membrane resistance $R$ and capacitance $C$.
@@
 In our framework, we adopt a backpropagation-based training method[8891809] for SNN.
-Since the activation function of the neuron in 10 is a discontinuous function, directly computing its derivative can lead to network instability.
+Since the activation function of the neuron in equation (10) is a discontinuous function, directly computing its derivative can lead to network instability.
 Consequently, various surrogate gradient methods have been proposed [8891809].
-A common approach is to use the original step function $y_{i}=\Theta(a_{i})$ during the forward pass, while replacing its derivative with $\theta^{\prime}(a_{i})=\sigma^{\prime}(a_{i})$ instead of $\Theta^{\prime}(a_{i})$ during backpropagation, where $\sigma(a_{i})$ is referred to as the surrogate activation function.
-Typically, $\sigma(a_{i})$ is a smooth and continuous function that resembles the shape of $\Theta(a_{i})$ while enabling stable gradient computation.
+A common approach is to use the original step function $y=\Theta(a)$ during the forward pass, while replacing its derivative with $\theta^{\prime}(a)=\sigma^{\prime}(a)$ instead of $\Theta^{\prime}(a)$ during backpropagation, where $\sigma(a)$ is referred to as the surrogate activation function.
+Typically, $\sigma(a)$ is a smooth and continuous function that resembles the shape of $\Theta(a)$ while enabling stable gradient computation.
 As illustrated in Fig.1c, the framework employs three types of neurons across all modules, represented in gray, yellow, and green.
@@
 In the machine learning domain, the task of fitting a curve to observed data under supervision is referred to as a regression task, for which extensive research and mature methods already exist[henkes2024spiking, robotics13090126].
-Figure 2: (a) Structure of one network unit during training in the cerebellum module. (b) Structure of cerebellum network during prediction.
+Figure 2: (a) Structure of one network unit in the cerebellum module during training. (b) Structure of cerebellum network during prediction.
 As shown in Fig.2a, to implement the cerebellum module, a four-layer recurrent SNN is designed as a network unit of the module, analogous to the “microcomplex” proposed in [abadia2019robot, abadia2021cerebellar].
@@
 These trained networks are then combined to form the complete cerebellum module structure illustrated in Fig.
-2b, enabling adaptive gravity compensation across multiple load conditions.
-In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$.
+2b.
+In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$.
 This matrix corresponds to the learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy conditions.
@@
 Figure 3: Structure of thalamus module.
-The thalamus module is implemented as a three-layer fully connected SNN, as illustrated in Fig.
+The thalamus module contains five layers of neurons, as illustrated in Fig.
 3.
-Its input, $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$, corresponds to the six-dimensional force/torque sensor data measured at the end-effector, while its output, $\bm{w}=(w_{1},w_{2})^{T}\in\mathbb{R}^{2}$, represents the weight coefficients assigned to the two gravity compensation torques provided by the cerebellum module.
+Its input $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$ and $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\in\mathbb{R}^{7\times 2}$, corresponds to the six-dimensional force/torque sensor data measured at the end-effector, and the cerebellum module output, respectively.
+Its output, $\bm{w}=(w_{1},w_{2})^{T}\in\mathbb{R}^{2}$, represents the weight coefficients assigned to the two gravity compensation torques provided by the cerebellum module.
 The input layer is fully connected to a hidden layer consisting of five LIF neurons, which in turn connects to an output layer of two non-spiking LIF neurons.
@@
 Following this encoding process, two sets of discrete spike trains are generated—one representing position information and the other representing velocity information.
-These spike trains are then transmitted to stimulate the activity of non-spiking LIF neurons in the output layer, with stimulation strength modulated by the feedback control gains $k_{p,i}$ and $k_{v,i}$ provided by the brainstem module, as illustrated by the blue and red dashed lines, respectively.
+These spike trains are then transmitted to stimulate the activity of non-spiking LIF neurons in the output layer, with stimulation strength modulated by the feedback control gains $k_{p,i}$ and $k_{v,i}$, $i=1,\cdots,7$ provided by the brainstem module, as illustrated by the blue and red dashed lines, respectively.
 Accordingly, at each time step, the membrane potential increments of the output non-spiking LIF neurons can be expressed as
@@
 The controlled robotic arm consists of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector.
+The control cycle of each layer is set as $T_{1}=1$ms, $T_{2}=10$ms, $T_{3}=50$ms.
 The simulation time step is set to 1 ms, and the control frequency is configured to 1 kHz.
@@
 while the average RMSE over an entire trajectory is defined as
-$\text{RMSE}=\frac{1}{N}\sum_{t=1}^{N}RMSE(t)$ (27)
+$\text{RMSE}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}(t)$ (27)
 Three types of trajectories are employed in the simulations: horizontal circle, inclined circle, and spatial figure-eight.
@@
 The temporal separation of these peaks validates the layered organization of the control framework.
-To verify this correspondence, ablation studies were conducted by selectively disabling individual modules, as shown in Fig.
+To verify this correspondence, ablation experiments were conducted by selectively disabling individual modules, as shown in Fig.
 6e, 6f and 6g.
@@
 In summary, the proposed framework exhibits robust adaptability and rapid convergence to load variations, with distinct loop-specific responses emerging in a biologically consistent sequence.
-This layered activation mirrors the progression of human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [forgaard2015voluntary], thereby reinforcing the plausibility and effectiveness of the neuromorphic design.
+This step-by-step activation phenomenon mirrors the progression of human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [forgaard2015voluntary], thereby reinforcing the plausibility and effectiveness of the neuromorphic design.
 V Experiment results
@@
 At each time step, the position and orientation RMSE are defined as
-$RMSE_{p}(t)=$ (29)
+$\text{RMSE}_{p}(t)=$ (29)
 $\sqrt{(x_{d}(t)-x(t))^{2}+(y_{d}(t)-y(t))^{2}+(z_{d}(t)-z(t))^{2}}$
-$RMSE_{o}(t)=$ (30)
+$\text{RMSE}_{o}(t)=$ (30)
 $\sqrt{(\theta_{d}(t)-\theta(t))^{2}+(\psi_{d}(t)-\psi(t))^{2}+(\phi_{d}(t)-\phi(t))^{2}}$
@@
 The average position and orientation RMSE of a trajectory are defined as
-$RMSE_{p}=\frac{1}{N}\sum_{t=1}^{N}RMSE_{p}(t)$ (31)
-$RMSE_{o}=\frac{1}{N}\sum_{t=1}^{N}RMSE_{o}(t)$ (32)
+$\text{RMSE}_{p}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{p}(t)$ (31)
+$\text{RMSE}_{o}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{o}(t)$ (32)
 where $N$ represents the trajectory length.
@@
 In this paper, we propose a CNS-inspired control framework based on SNN, enabling agile robotic arm control in complex environments.
-The framework introduces a hierarchical architecture with five modules, is fully implemented using SNNs, and has been validated through both simulations and real-world robotic experiments, outperforming an industrial-grade baseline.
+The framework introduces a hierarchical architecture with five modules implemented using SNNs, and has been validated through both simulations and real-world robotic experiments.
+Results show that this framework operates in a manner similar to the human CNS and can outperform an industrial-grade baseline.
 Future work will focus on improving reinforcement learning for higher reproducibility, enhancing the cerebellar module for better gravity compensation, and further increasing control accuracy.
@@
 Parameter Value Parameter Value Parameter Value
-$(x_{0},y_{0},z_{0})$
-(0.54,0.0,
-0.45) m
-$R_{0}$ 0.14 m $T_{0}$ 4 s
-$(x_{1},y_{1},z_{1})$
-(0.6,0.35,
-0.0) m
-$R_{1}$ 0.2 m $T_{1}$ 4 s
-$(x_{2},y_{2},z_{2})$
-(0.63,-0.11,
-0.3) m
-$R_{2}$ 0.14 m $T_{2}$ 4 s
-$(x_{3},y_{3},z_{3})$
-(0.0,0.61,
-0.3) m
-$R_{3}$ 0.14 m $T_{3}$ 4s
-$(x_{4},y_{4},z_{4})$
-(0.53,-0.11,
-0.4) m
-$R_{4}$ 0.2 m $T_{4}$ 3s
-$(x_{5},y_{5},z_{5})$
-(0.53,-0.11,
-0.4) m
-$R_{5}$ 0.2 m $T_{5}$ 2s
+$(x_{0},y_{0},z_{0})$ (0.54,0.0, 0.45) m $R_{0}$ 0.14 m $T_{0}$ 4 s
+$(x_{1},y_{1},z_{1})$ (0.6,0.35, 0.0) m $R_{1}$ 0.2 m $T_{1}$ 4 s
+$(x_{2},y_{2},z_{2})$ (0.63,-0.11, 0.3) m $R_{2}$ 0.14 m $T_{2}$ 4 s
+$(x_{3},y_{3},z_{3})$ (0.0,0.61, 0.3) m $R_{3}$ 0.14 m $T_{3}$ 4s
+$(x_{4},y_{4},z_{4})$ (0.53,-0.11, 0.4) m $R_{4}$ 0.2 m $T_{4}$ 3s
+$(x_{5},y_{5},z_{5})$ (0.53,-0.11, 0.4) m $R_{5}$ 0.2 m $T_{5}$ 2s
 $\theta_{1}$ $\frac{\pi}{6}$ rad $\phi_{1}$ $\frac{\pi}{3}$ rad $T_{6}$ 3s
 $\theta_{2}$ $\frac{\pi}{6}$ rad $T_{7}$ 3s
-[Uncaptioned image] Yanbo Pang received the B.Eng.
+Yanbo Pang received the B.Eng.
 from Sun Yat-sen University in 2022 and the M.Sc.
 from Tsinghua University in 2025.
-He is currently a Ph.D student at the Technical University of Munich (TUM).
+He is currently an intern at Booster Robotics.
 His research interests include neuromorphic computing, neurorobotics and robotic arm control.
-[…]
-[…]
+Qingkai Li received his B.S.
+from Tsinghua University in 2021.
+He is currently a Ph.D.
+candidate in the Department of Automation at Tsinghua University, with research interests spanning neuromorphic computing, bio-inspired algorithms, imitation learning, and whole-body control.
+Mingguo Zhao received the B.Eng., M.Sc.
+and Ph.D.
+from Harbin Institute of Technology in 1995, 1997 and 2001, respectively.
+He is currently a Full Professor with the Department of Automation and a director of the Beijing Innovation Center for Future Chips, Tsinghua University.
+His research interests include brain-inspired robot control and general robotic platform based on neuromorphic computing and hardware.

### v3

@@
-CBMC-V3: A CNS-inspired Control Framework Towards Manipulation Agility with SNN
+CBMC-V3: A CNS-inspired Control Framework Towards Agile Manipulation with SNN
 Abstract
-As robotic arm applications extend beyond industrial settings into service-oriented sectors such as catering, household and retail, existing control algorithms struggle to achieve the agile manipulation required for complex environments with dynamic trajectories, unpredictable interactions, and diverse objects.
-This paper presents a biomimetic control framework based on Spiking Neural Networks (SNN), inspired by the human Central Nervous System (CNS), to achieve agile control in such environments.
-The proposed framework features five control modules (cerebral cortex, cerebellum, thalamus, brainstem, and spinal cord), three hierarchical control levels (first-order, second-order, and third-order), and two information pathways (ascending and descending).
-Each module is fully implemented using SNN.
-The spinal cord module uses spike encoding and Leaky Integrate-and-Fire (LIF) neurons for feedback control.
-The brainstem module employs a network of LIF and non-spiking LIF neurons to dynamically adjust spinal cord parameters via reinforcement learning.
-The thalamus module similarly employs a network of LIF and non-spiking LIF neurons to adjusts the cerebellum’s torque outputs via reinforcement learning.
-The cerebellum module, which provides feedforward gravity compensation torques, uses a recurrent SNN to learn the robotic arm’s dynamics through regression.
-The framework is validated both in simulation and on real-world robotic arm platform under various loads and trajectories.
-Results demonstrate that our method outperforms the industrial-grade position control in manipulation agility.
+As robotic arm applications expand beyond traditional industrial settings into service-oriented domains such as catering, household, and retail, existing control algorithms struggle to achieve the level of agile manipulation required in unstructured environments characterized by dynamic trajectories, unpredictable interactions, and diverse objects.
+This paper presents a biomimetic control framework based on spiking neural network (SNN), inspired by the human central nervous system (CNS), to address these challenges.
+The proposed framework comprises five control modules—cerebral cortex, cerebellum, thalamus, brainstem, and spinal cord—organized into three hierarchical control levels (first-order, second-order, and third-order) and two information pathways (ascending and descending).
+All modules are fully implemented using SNN.
+The framework is validated through both simulation and experiments on a commercial robotic arm platform across a range of control tasks.
+The results demonstrate that the proposed method outperforms the baseline in terms of agile motion control capability, offering a practical and effective solution for achieving agile manipulation.
 I Introduction
-In recent years, robotic arms are increasingly applied in service-oriented sectors such as catering, household and retail.
-Compared with traditional industrial settings, these scenarios feature more complex environments and diverse tasks while requiring better agility, thereby posing new challenges to control strategies.
-For instance, a 7-DOF robotic arm can readily cut a circular contour on a workpiece or transport heavy replacement parts, yet it still struggles to perform a task as subtle as peeling an apple, which can be easily accomplished by the human arm.
-Compared with biological systems, even the most advanced robotic arms today fall short of the agility of human upper limb [mason2018], underscoring the existence of advanced motor control mechanisms in humans that remain underutilized in robotic systems.
-In well-modeled environments, control algorithms combining Whole-Body Control (WBC) [Abe2007, liu2011interactive, weighttans2011, Lee2012Intermediate, Escande2014, Kim2019, han2021recursive] and Model Predictive Control (MPC) [carron2019data, pankert2020perceptive, wang2023trajectory] can effectively counteract disturbances and achieve strong control performance.
-Such model-based approaches reflect human capabilities in applying prior knowledge to solve specific tasks.
-However, they lack the adaptability inherent to human motor learning.
-In complex scenarios, variations in the size, mass, or materials of manipulated objects may render the kinematic and dynamic models inaccurate, limiting the effectiveness of these model-based strategies.
-With the rapid progress of neural networks, their powerful approximation and learning capabilities have motivated researchers to employ them for adaptive model correction in response to environmental changes.
-For example, [chaoui2009ann] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [he2017adaptive] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [wang2018adaptive] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [salloom2020adaptive] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [liu2019adaptive] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [pham2020adaptive] applied radial basis function networks with dynamic surface control for robust dual-arm coordination.
-Despite these progression, they rely on artificial neural networks (ANN), which inherently disregard the temporal-spatial characteristics of motion, focusing only on instantaneous control.
-Furthermore, they remain constrained by robotic dynamics models, with network structures and training methods often tailored to specific problems and tightly coupled to complex mathematical formulations, thereby limiting generalizability.
-Recently, spiking neural networks (SNN) combined with brain-inspired architectures have emerged as promising alternatives.
-By emulating biological neurons, SNNs process and store information through spike trains and membrane potential dynamics, offering advantages in computational speed, energy efficiency, information dimension, and biological plausibility, making them well-suited for robotic control [bing2018survey].
-For instance, [carrillo2008real] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn manipulator dynamics for torque control; [abadia2019robot] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [abadia2021cerebellar] further established resilience to control delays, highlighting potential in teleoperation and cloud robotics; [doi:10.1126/scirobotics.adp2356] refines this framework to enable robotic arms to dynamically adjust compliance.
-However, these methods suffer from structural complexity, computational cost, hyperparameter sensitivity, and limited interpretability, as they focus narrowly on cerebellar functionality while neglecting the integrative role of the broader central nervous system (CNS), which coordinates control through interactions among several neural areas[Halassa_2022].
-Inspired by this systemic perspective of the CNS, [li2023cbmc] proposed the CBMC framework, incorporating modules such as the cortex, cerebellum, and spinal cord.
-Yet its computational inefficiency restricted real-time applicability.
-Building on this, [10769842] introduced a hierarchical control architecture deployable on neuromorphic hardware, though its lack of neural networks in some modules compromised accuracy.
-Despite these advances, current brain-inspired algorithms have yet to demonstrate agile performance comparable to that of humans.
-As defined in [8610238], agility is a capability characterized by rapid and efficient transitions between diverse motion patterns, with precision and speed as its key attributes.
-In this work, we propose CBMC-V3, a fully SNN-based brain-inspired control framework designed to achieve agile robotic arm control.
-The key contributions are fourfold: (1) Proposing a CNS-inspired control framework leveraging hierarchical modularity; (2) Designing a fully SNN-based implementation; (3) Validating the framework and module design in simulation experiments; (4) Deploying the algorithm on a robotic platform, demonstrating agility.
-The remainder of this paper is organized as follows: Section II introduces the framework structure; Section III details the SNN-based implementations of individual modules; Section IV presents simulation analyses; Section V conducts experiments on a robotic arm platform, and Section VI concludes the study.
-Figure 1: Overview of the human CNS and the proposed framework structure.
-(a) Human CNS relating to motion control.
-Sensory information and motor commands are transmitted through three orders of neurons along ascending and descending pathways.
-Sensory signals travel from peripheral receptors to the spinal cord, brainstem, and ultimately the cerebellum and cerebral cortex, while motor commands descend from the cerebral cortex and cerebellum through the thalamus and spinal cord to the muscles.
-The spinal cord can independently generate reflexive responses via reflex arcs, enabling rapid reactions to stimuli [LLOBERA2023190].
-The brainstem contributes to movement termination, reorientation, and coordination with the spinal cord [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137].
-The thalamus functions as a hub for motor coordination, multimodal perception, and sensorimotor integration [LLOBERA2023190, Halassa_2022].
-The cerebellum regulates body movements, balance, and posture by integrating sensory inputs and refining motor commands [LLOBERA2023190].
-The cerebral cortex generates decisions and motion trajectories.
-(b) Schematic diagram of our proposed framework, featuring five modules, three control levels and two pathways as mentioned in (a).
-(c) Neural network design of each module, utilizing three types of spiking neurons: input fiber (grey), LIF neuron (yellow) and non-spiking LIF neuron (green).
+In recent years, robotic arms have been increasingly deployed in manipulation-intensive domains such as catering, household, and retail services, posing new challenges for control strategies.
+While a 7-DOF robotic arm can readily cut precise circular contours or transport heavy replacement parts, it still struggles with delicate tasks such as peeling an apple—an operation that the human arm performs effortlessly.
+This gap largely arises from humans’ strong motion agility, which enables robust and flexible handling of delicate, contact-rich interactions.
+However, defining and quantifying agility in motion control remains challenging due to its context-dependent and multifaceted nature.
+Drawing insights from [39, 12], we characterize agile motion control using four key features: (1) precision, (2) adaptation, (3) resilience, and (4) energy efficiency.
+Precision measures an algorithm’s ability to accurately track a desired trajectory, for example in terms of end-effector position and orientation tracking errors.
+Adaptation evaluates the ability to maintain tracking precision across different tasks, such as under varying load masses and trajectory configurations.
+Resilience reflects how quickly the algorithm recovers from unstructured interactions or disturbances.
+Energy efficiency quantifies the computational resources required by the algorithm.
+Control algorithms that combine Whole-Body Control (WBC) [4, 27, 36, 24, 13, 23, 19] and Model Predictive Control (MPC) [9, 33, 44] can achieve high precision in well-defined environments and typically exhibit high energy efficiency.
+However, their resilience is often compromised as model complexity increases.
+Moreover, these methods lack effective mechanisms to cope with environmental variations and model inaccuracies, such as objects with differing shapes or masses and elastic or biomimetic joints, resulting in poor adaptation.
+Artificial neural network (ANN) can learn and adapt to environmental changes, resulting in relatively strong adaptation capabilities.
+When combined with model-based dynamics, it can also achieve relatively high precision.
+For example, [10] designed feedforward and feedback neural networks to compensate for joint flexibility and friction; [20] employed fuzzy neural networks with adaptive impedance control to handle unknown dynamics and state constraints; [43] proposed a neural-network-based adaptive controller capable of addressing nonlinear dynamics and external disturbances without prior models; [37] integrated genetic algorithm observers with neural networks to compensate for underwater disturbances; [26] investigated the role of hidden-layer size in adaptive control and proposed dynamic adjustment strategies; and [35] applied radial basis function networks with dynamic surface control for robust dual-arm coordination.
+However, these methods typically exhibit limited resilience and energy efficiency due to the computational characteristics of ANN [6].
+Moreover, the limited ability of ANN to model temporal dynamics further restricts their adaptation in highly dynamic manipulation tasks.
+Recently, spiking neural network (SNN) combined with cerebellum-inspired architecture has emerged as promising alternatives.
+By emulating biological neurons, SNN processes and stores information through spike trains and membrane potential dynamics, offering significant advantages in energy efficiency [6].
+[8] developed a cerebellum-inspired SNN trained via spike-timing-dependent plasticity (STDP) to learn robotic arm dynamics for torque control; [2] extended this model to human–robot collaboration scenarios, demonstrating robustness under unknown disturbances; [3] addressed control delays using this network, highlighting its potential for teleoperation and cloud robotics; and [1] further refined the framework to enable robotic arms to dynamically adjust compliance.
+By modeling human neural processing at the level of individual neurons and closely mimicking the structure of the human cerebellum, these approaches can exhibit human-like behavior and strong adaptation capabilities.
+However, despite the low per-neuron energy consumption, the extremely large network scale (e.g., over 60k neurons and 32M synaptic connections) limits the overall energy efficiency of the framework.
+Moreover, the use of unsupervised, biologically inspired learning rules such as STDP—whose convergence properties lack theoretical guarantees—constrains achievable precision and resilience.
+In our previous work CBMC-V2 [32], we presented a hierarchical control framework inspired by the central nervous system (CNS).
+Compared with cerebellum-inspired SNN approaches, this framework achieves higher energy efficiency by modular design and neuromorphic hardware deployment, and attains improved precision and resilience by incorporating gradient-based learning into both the cerebral cortex and cerebellum modules.
+However, the framework exhibits weak adaptation due to limited learning capability, arising from insufficient sensory inputs (e.g., load information), the absence of SNN in some modules, and the tight coupling between control and planning in the cerebral cortex module.
+Furthermore, because the cerebellum module largely preserves the network structure adopted in [2, 3, 1], the framework inherits their inherent limitations, which constrain the achievable precision, resilience, and energy efficiency.
+Overall, this research aims to provide a motion control framework towards agile manipulation by addressing the limitations identified in CBMC-V2.
+To this end, we propose CBMC-V3, a CNS-inspired, fully SNN-based control framework.
+The main contributions of this work are threefold: (1) the proposal of a CNS-inspired control framework; (2) the design of a fully SNN-based implementation; and (3) the demonstration of agile motion control in both simulation and on a commercial robotic arm.
+The remainder of this paper is organized as follows: Section II introduces the framework structure; Section III details the SNN-based implementations of individual modules; Section IV presents simulation results; Section V reports experiment results on a commercial robotic arm, and Section VI concludes this paper.
+Figure 1: Overview of the human CNS and the proposed framework.
+(a) Human CNS relating to motor control.
+(b) Diagram of our proposed framework, featuring five modules, three control levels and two pathways.
+(c) Neural network design of each module with three types of spiking neurons: connecting fiber (grey), LIF neuron (yellow) and non-spiking LIF neuron (green).
 II Framework structure
-Inspired by the motion control mechanism of the human CNS, as illustrated in Fig.1a, the structure of our proposed framework CBMC-V3 is shown in Fig.1b.
-This framework consists of three fundamental components: (1) five functional modules, namely the cerebral cortex module, cerebellum module, thalamus module, brainstem module, and spinal cord module; (2) three hierarchical control levels, namely the first-order, second-order, and third-order level; (3) two information pathways, the ascending and descending pathway.
-In the following sub-sections, we will introduce each of these components in detail and explain how the framework integrates functional modules, control loops, and information pathways to accomplish a complete operational cycle.
-II-A Module function
-In the human CNS, the spinal cord can generate reflexive actions independently via reflex arcs, enabling rapid responses to sensory stimuli without cortical input [LLOBERA2023190].
-Inspired by this mechanism, the spinal cord module in our framework implements a PD-like feedback control function, producing torque commands at real-time control frequencies, i.e.,
+Inspired by the motor control mechanism of the human CNS, as illustrated in Fig.1a, the structure of our proposed framework CBMC-V3 is shown in Fig.1b.
+This framework consists of three key aspects: (1) five modules including the cerebral cortex module, the cerebellum module, the thalamus module, the brainstem module, and the spinal cord module; (2) three hierarchical control levels, namely the first-order level, second-order level, and third-order level; (3) two information pathways, i.e.
+the ascending pathway and the descending pathway.
+In the following sub-sections, we will introduce each of these aspects and present the pseudocode implementation of this framework.
+II-A Modules
+II-A1 The cerebral cortex module
+The cerebral cortex module makes decisions and generates motion trajectories.
+In this research, it stores and replays pre-planned velocity $\bm{\dot{q}}_{d}\in\mathbb{R}^{n}$ and position $\bm{q}_{d}\in\mathbb{R}^{n}$ of the $n$ actuated joints.
+II-A2 The cerebellum module
+The cerebellum module is a cerebellar-like network pre-trained to store gravity compensation motor patterns for corresponding load masses.
+In human motor control, the cerebellum is responsible for coordinating and regulating body movements, ensuring muscle and postural balance.
+It integrates sensory inputs, such as muscle tension and joint position, and adjusts the magnitude and direction of motor commands to achieve precise and fluid motion [28].
+The gravity compensation torque for a robotic arm is designed to counteract the gravitational forces acting on the arm itself.
+It plays a crucial role in ensuring the smoothness and stability of motion.
+The cerebellum can be regarded as performing a function similar to gravity compensation, and the cerebellum module in our framework is therefore designed to provide this compensation.
+Furthermore, we notice that during human manipulation, accurate dynamic model is not required.
+Instead, the brain relies on stored motor patterns to cope with different load conditions.
+For instance, when grasping an object, we estimate its mass based on prior experience and adjust the exerted force accordingly.
+Such knowledge is typically not acquired on site during execution, but rather accumulated over time through long-term daily interactions.
+Inspired by this biological mechanism, we equip the cerebellum module with the ability to pre-learn $m$ gravity compensation patterns corresponding to different load masses.
+These patterns can then be retrieved and combined in prediction to generate a compensation torque matrix
+$\begin{split}\bm{T}_{{\text{cb}}}(\bm{q})&=(\bm{\tau}_{\text{cb},1}(\bm{q}),\bm{\tau}_{\text{cb},2}(\bm{q}),\cdots,\bm{\tau}_{\text{cb},m}(\bm{q}))\\ &=\begin{pmatrix}\tau_{\text{cb},11}&\cdots&\tau_{\text{cb},1m}\\ \vdots&\ddots&\vdots\\ \tau_{\text{cb},n1}&\cdots&\tau_{\text{cb},nm}\end{pmatrix}\end{split}$ (1)
+as a function of actual joint positions $\bm{q}\in\mathbb{R}^{n}$, where $\tau_{\text{cb},ji}(\bm{q}),j=1,\cdots,n,i=1,\cdots,m$ denotes the torque provided for the $j$-th joint under the $i$-th compensation pattern.
+II-A3 The thalamus module
+The thalamus module learns to generate compensation torque adaptable to arbitrary loads by online RL using proprioceptive information and the cerebellum module output.
+In unstructured environments, the selection of motor patterns may be inaccurate due to factors such as visual deception.
+For instance, the mass of an opaque box cannot be reliably estimated visually, which may lead to insufficient force to move the load or, conversely, an excessive initial force causing large errors.
+Nevertheless, humans can quickly adjust motor output upon contacting the object, using sensory feedback to correct their movement patterns.
+This phenomenon is closely associated with the function of the thalamus, which serves as a central hub for motor coordination and plays a critical role in multimodal perception and sensorimotor integration [28, 18].
+Analogously, a robotic arm can perceive the load mass via force sensors installed at the end-effector.
+The thalamus module, similar to humans, uses the sensed load information to appropriately integrate the torque matrix provided by the cerebellum module, i.e.,
+$\begin{split}\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})&=\bm{T}_{\text{cb}}(\bm{q})\cdot\bm{w}(\bm{F}_{\text{ef}})\\ &=\begin{pmatrix}\tau_{g,1}\\ \vdots\\ \tau_{g,n}\end{pmatrix}=\begin{pmatrix}w_{1}\tau_{\text{cb},11}+\cdots+w_{m}\tau_{\text{cb},1m}\\ \vdots\\ w_{1}\tau_{\text{cb},n1}+\cdots+w_{m}\tau_{\text{cb},nm}\end{pmatrix}\end{split}$ (2)
+where $\bm{F}_{\text{ef}}$ denotes the end-effector torque sensor feedback, $\bm{w}(\bm{F}_{\text{ef}})=(w_{1},w_{2},\dots,w_{m})^{\top}$ represents the $m$ dynamically learned weight coefficients of the thalamus module, and $\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{g,1},\dots,\tau_{g,n})^{\top}$ are the resulting gravity compensation torque of each joint for arbitrary load mass.
+By combining the prior knowledge stored in the cerebellum module with the adaptive weighting capability of the thalamus module, the framework effectively addresses the unknown dynamics of the robotic arm’s end-effector load.
+II-A4 The brainstem module
+The brainstem module learns to provide appropriate feedback control coefficients by online RL using proprioceptive information.
+The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [29, 25].
+Motivated by this, the brainstem module in our framework dynamically adjusts the proportional and derivative coefficients of the $n$ actuated joints based on trajectory information, i.e.,
+$diag(k_{p,1},k_{p,2},\cdots,k_{p,n})=\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ (3)
+$diag(k_{v,1},k_{v,2},\cdots,k_{v,n})=\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ (4)
+where $\bm{\dot{q}}\in\mathbb{R}^{n}$ is the actual velocity of the $n$ actuated joints.
+II-A5 The spinal cord module
+The spinal cord module combines feedback control coefficients from the brainstem module and the compensation torque from the thalamus module to generate real-time control torque commands.
+In the human CNS, the spinal cord can generate reflexive actions independently via reflex arcs, enabling rapid responses to sensory stimuli without cortical input [28].
+Inspired by this mechanism, the spinal cord module produces torque commands at real-time control frequency with a PD-like feedback control mechanism, i.e.,
 $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ $=\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})(\bm{\dot{q}_{d}}-\bm{\dot{q}})$ (5)
 $+\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})(\bm{q_{d}}-\bm{q})$
-The brainstem has been shown to contain pathways involved in movement termination and reorientation, and to interact with the spinal cord to coordinate motor behaviors [merel2019hierarchical, annurev:/content/journals/10.1146/annurev-neuro-082321-025137].
-Motivated by this, the brainstem module in our framework dynamically adjusts the PD gains for each joint based on trajectory information, providing the spinal cord module with time-varying proportional and derivative coefficients, i.e.,
-$\bm{K}_{V}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{v,1},k_{v,2},\cdots,k_{v,n})$ (2)
-$\bm{K}_{P}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})=diag(k_{p,1},k_{p,2},\cdots,k_{p,n})$ (3)
-By integrating spinal feedback control with brainstem-driven adaptive gain modulation, the framework circumvents the complexity of manipulator dynamics while compensating for structured uncertainties through learning.
-The gravity compensation torque of a robotic manipulator, denoted as $\bm{\tau}_{g}(\bm{q})$, is a feedforward torque designed to counteract the gravitational forces acting on the manipulator itself.
-It plays a crucial role in ensuring the smoothness and stability of motion.
-In human motor control, the cerebellum is primarily responsible for coordinating and regulating body movements, ensuring muscle and postural balance.
-It integrates sensory inputs, such as muscle tension and joint position, and adjusts the magnitude and direction of motor commands to achieve precise and fluid motion [LLOBERA2023190].
-Analogously, gravity compensation in robotic manipulators serves a function comparable to that of the cerebellum, and thus the cerebellar module in our framework is tasked with providing this compensation.
-Furthermore, we note that during human arm manipulation, accurate models of dynamics are not strictly required.
-Instead, the brain relies on stored motor patterns to cope with different loading conditions.
-For instance, when grasping an object, we estimate its weight based on prior experience and adjust the exerted force accordingly.
-Such experiential knowledge is typically not acquired on site during execution, but rather accumulated over time through long-term daily interactions.
-Inspired by this biological mechanism, the proposed framework equips the cerebellum module with the ability to pre-learn multiple gravity compensation ‘patterns’ corresponding to different load masses.
-These patterns can then be retrieved and combined to generate an $n\times m$ set of compensation torques,
-$\begin{split}\bm{T}_{{\text{cb}}}(\bm{q})&=(\bm{\tau}_{\text{cb},1}(\bm{q}),\bm{\tau}_{\text{cb},2}(\bm{q}),\cdots,\bm{\tau}_{\text{cb},m}(\bm{q}))\\ &=\begin{pmatrix}\tau_{\text{cb},11}&\cdots&\tau_{\text{cb},1m}\\ \vdots&\ddots&\vdots\\ \tau_{\text{cb},n1}&\cdots&\tau_{\text{cb},nm}\end{pmatrix}\end{split}$ (4)
-as a function of joint positions, where $\tau_{\text{cb},ji}(\bm{q}),j=1,\cdots,n,i=1,\cdots,m$, denotes the torque provided for the $j$-th joint under the $i$-th compensation pattern.
-Although suitable motor patterns can often be selected based on prior experience, in certain cases, such selections may be inaccurate due to factors such as visual deception.
-For instance, the mass of an opaque box cannot be reliably estimated visually, which may lead to insufficient force to move the load or, conversely, an excessive initial force causing large errors.
-Nevertheless, humans can quickly adjust their motor output upon physically contacting the object, using sensory feedback to correct their movement patterns.
-This phenomenon is closely associated with the function of the thalamus, which serves as a central hub for motor coordination and plays a critical role in multimodal perception and sensorimotor integration [LLOBERA2023190, Halassa_2022].
-Analogously, a robotic arm can ‘perceive’ the load weight via force sensors installed at the end-effector.
-The thalamus module, similar to humans, uses the sensed load information to dynamically weight and integrate the torque contributions provided by the cerebellar module, generating an actual gravity compensation torque, i.e.,
-$\begin{split}\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})&=\bm{T}_{\text{cb}}(\bm{q})\cdot\bm{w}(\bm{F}_{\text{ef}})\\ &=\begin{pmatrix}\tau_{g,1}\\ \vdots\\ \tau_{g,n}\end{pmatrix}=\begin{pmatrix}w_{1}\tau_{\text{cb},11}+\cdots+w_{m}\tau_{\text{cb},1m}\\ \vdots\\ w_{1}\tau_{\text{cb},n1}+\cdots+w_{m}\tau_{\text{cb},nm}\end{pmatrix}\end{split}$ (5)
-where $\bm{F}_{\text{ef}}$ denotes the end-effector torque sensor feedback, $\bm{w}(\bm{F}_{\text{ef}})=(w_{1},w_{2},\dots,w_{m})^{T}$ represents the dynamically learned weight coefficients of the thalamus module, and $\tau_{g,1},\dots,\tau_{g,n}$ are the resulting gravity compensation torques.
-By combining the prior knowledge stored in the cerebellum module with the adaptive weighting capability of the thalamus module, the framework effectively addresses the unknown dynamics of the manipulator’s end-effector load.
-The resulting torque is then summed with the feedback torque generated by the spinal cord module and applied as joint-level control commands to the actuators, i.e.,
-$\bm{\tau}=\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})+\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{1},\cdots,\tau_{n})^{T}$ (6)
-where $\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})$ is produced by the spinal cord and brainstem module, $\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})$ is produced by the cerebellum and thalamus module, and $(\tau_{1},\dots,\tau_{n})^{T}$ represents the joint torques delivered to the actuators.
-In conclusion, the function of each module can be summarized as follows:
-  • The cerebral cortex module makes decisions and generates motion trajectories. In this research, it stores and replays pre-planned trajectories.
-  • The cerebellum module is a cerebellar-like network pre-trained to store several fundamental gravity compensation motion patterns.
-  • The thalamus module learns to generate compensation torque adaptable to varying loads by online RL using proprioception information and the cerebellum module output.
-  • The brainstem module learns to provide appropriate feedback control coefficients by online RL using proprioception information.
-  • The spinal cord module combines feedback control coefficients from the brainstem and torque from the thalamus module to generate real-time control torque commands.
-II-B Hierarchical control level
-Clearly, different neural regions in the central nervous system operate hierarchically during motion control, with each level specializing in particular tasks while communicating and cooperating through neural pathways.
-This hierarchical organization enables organisms to execute a wide range of motion tasks flexibly and stably across varying environments.
+By integrating spinal feedback control with brainstem-driven adaptive coefficients modulation, the framework effectively addresses non-linearity, Coriolis and centrifugal forces from unknown trajectories.
+This reflexive torque is then summed with the compensation torque from the thalamus to generate control torque command $\bm{\tau}$ for actuators, i.e.,
+$\bm{\tau}=\bm{\tau}_{\text{fb}}(\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q})+\bm{\tau}_{g}(\bm{q},\bm{F}_{\text{ef}})=(\tau_{1},\cdots,\tau_{n})^{\top}$ (6)
+II-B Control levels
+Different neural regions in the CNS operate hierarchically during motor control, with each level specializing in particular tasks while communicating and cooperating through neural pathways.
+This hierarchical organization enables organisms to execute a wide range of motion tasks flexibly and stably in unstructured environments.
 Physiological studies indicate that the multilevel transmission of neural signals is a key component of this hierarchy.
-Proprioceptive information is conveyed from sensory receptors to the brain via a three-order neuronal pathway[william_ascending_pathway, LLOBERA2023190], as illustrated in Fig.1a.
+Proprioceptive information is conveyed from sensory receptors to the brain via a three-order neuronal pathway[28, 30, 40], as illustrated in Fig.1a.
 First-order neurons receive impulses from the skin and muscles and transmit them to the spinal cord.
@@
   • First-order level: Comprising the spinal cord module, this level operates at the highest frequency.
-Its primary function is to generate real-time control torques sent to the robotic arm, ensuring accurate and rapid responses.
-  • Second-order level: Including the brainstem and thalamus module, this level operates at an intermediate frequency.
-It serves as a relay between the first-order and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order level and integrating and filtering the torques provided by the third-order level to supply feedforward torque to the first-order level.
-  • Third-order level: Comprising the cerebellum and cerebral cortex module, this layer operates at the lowest frequency.
-Its role is to generate movement trajectories and provide gravity compensation torques based on motion patterns, maintaining movement stability and smoothness.
-II-C Information pathway
-As illustrated in Fig.1a, the human central nervous system employs ascending and descending pathways for sensory information transmission and motor command execution, respectively.
-The ascending pathways primarily convey peripheral sensory information to the brain for perception and integration, while the descending pathways transmit motor commands from the brain to the spinal cord and peripheral nerves to control movements.
+It generates real-time control torques sent to the robotic arm, ensuring accurate and rapid responses.
+  • Second-order level: Including the brainstem module and thalamus module, this level operates at an intermediate frequency.
+It serves as a relay between the first-order and third-order levels, performing two main functions: regulating the feedback control coefficients of the first-order level, and integrating and filtering the torques provided by the third-order level to supply gravity compensation torque to the first-order level.
+  • Third-order level: Comprising the cerebellum module and cerebral cortex module, this layer operates at the lowest frequency.
+It generates motion trajectories and provide compensation torques based on motor patterns, maintaining movement stability and smoothness.
+II-C Information pathways
+As illustrated in Fig.1a, the human CNS employs ascending and descending pathways for sensory information transmission and motor command execution, respectively[40, 30].
+The ascending pathway conveys peripheral sensory information to the brain for perception and integration, while the descending pathway transmits motor commands from the brain to the spinal cord and peripheral nerves to control movements.
 Following this principle, our framework incorporates two analogous information pathways:
-  • Ascending pathway: Inspired by the dorsal column–medial lemniscal and spinocerebellar tracts[william_ascending_pathway], this pathway transmits proprioceptive and touch information, namely joint positions, velocities, and end-effector torque.
+  • Ascending pathway: This pathway transmits proprioceptive and touch information, namely joint positions, velocities, and end-effector torque feedback.
 It simulates the process by which muscle receptors relay signals to the CNS for processing and can be regarded as the input process of the framework.
-  • Descending pathway: Inspired by the corticospinal and reticulospinal tracts[william_ascending_pathway], this pathway transmits joint torque commands and parameter adjustment signals.
+  • Descending pathway: This pathway transmits joint torque commands and module coefficients.
 It simulates the process by which the CNS delivers motor commands to muscle effectors and can be regarded as the output process of the framework.
-II-D A complete operational cycle
-To clarify the interplay among modules, levels, and pathways, we illustrate in Fig.
-1b a complete working cycle of the proposed framework, showing how sensory inputs are transformed into motor control outputs.
-In each cycle, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$ and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$) are transmitted via first-order neurons to the spinal cord.
-Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem, which adaptively provides PD gains $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum, which outputs gravity compensation matrix $\bm{T}_{\text{cb}}$.
-These ascending transmissions complete the flow of proprioceptive information from sensors to the modules.
-On the descending side, $\bm{T}_{\text{cb}}$ is conveyed via third-order neurons to the thalamus, where end-effector torque feedback $\bm{F}_{\text{ef}}$ is integrated to generate a weight vector $\bm{w}$.
-This vector filters $\bm{T}_{\text{cb}}$ to produce the effective gravity compensation torque $\bm{\tau}_{g}$, which is passed via second-order neurons to the spinal cord.
-There, $\bm{\tau}_{g}$ is combined with $\bm{\tau}_{\text{fb}}$ to form the final control torque $\bm{\tau}$, which is delivered through first-order neurons to the actuators.
-Thus, the descending pathway completes motor command transmission from the modules to the joints.
-Because information reaches different modules through pathways of varying length, the framework is naturally organized into three control levels.
-The first-order level involves only spinal processing and yields the shortest latency and highest frequency.
-The second-order level adds brainstem and thalamic processing, resulting in longer pathways and lower frequency.
-The third-order level incorporates the cerebellum, producing the longest transmission path and the lowest frequency.
-The pseudocode implementation of this framework is provided in Algorithm 1.
+II-D Pseudocode implementation
+To clarify the interplay among modules, levels, and pathways illustrated in Fig. 1b, the pseudocode implementation of this framework is provided in Algorithm 1.
+Information travels different length to reach different modules along the pathway, so the framework is organized into three control levels with different frequencies.
+On the ascending side, proprioceptive data (actual joint states $(\bm{\dot{q}},\bm{q})$ and desired joint states $(\bm{\dot{q}}_{d},\bm{q}_{d})$) are transmitted via first-order neurons to the spinal cord module.
+Part of this input is locally processed to generate feedback torque $\bm{\tau}_{\text{fb}}$; part is relayed via second-order neurons to the brainstem module, which adaptively provides feedback coefficients $(\bm{K}_{P},\bm{K}_{V})$; and part is sent via third-order neurons to the cerebellum module, which outputs compensation matrix $\bm{T}_{\text{cb}}$.
+On the descending side, $\bm{T}_{\text{cb}}$ is conveyed via third-order neurons to the thalamus module.
+It uses end-effector torque feedback $\bm{F}_{\text{ef}}$ to generate a weight vector $\bm{w}$ which filters $\bm{T}_{\text{cb}}$ to produce the gravity compensation torque $\bm{\tau}_{g}$.
+This torque is passed via second-order neurons to the spinal cord module.
+In the spinal cord module, $\bm{\tau}_{g}$ is combined with $\bm{\tau}_{\text{fb}}$ to form the control torque command $\bm{\tau}$, which is delivered through first-order neurons to the actuators.
 Algorithm 1 CBMC-V3
-0:  joint proprioceptive sensor data $\bm{\dot{q}},\bm{q}$, planned trajectory data $\bm{\dot{q}}_{d},\bm{q}_{d}$, end-effector torque sensor data $\bm{F}_{ef}$
+0:  actual joint velocity and position $\bm{\dot{q}},\bm{q}$, planned joint velocity and position $\bm{\dot{q}}_{d},\bm{q}_{d}$, end-effector torque sensor data $\bm{F}_{ef}$
 0:  joint control torque $\bm{\tau}$
@@
 5:  initialize control torque variable $\bm{\tau}$
-6:  initialize intermediate variables $\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g},\bm{w},\bm{T}_{{\text{cb}}}$
-7:  read trajectory length $M$
-8:  for $t=1$ to $M$ do
+6:  initialize intermediate variables $\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g},\bm{T}_{{\text{cb}}}$
+7:  read trajectory length $L$
+8:  for $t=1$ to $L$ do
 9:   read data from sensors and update $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q},\bm{F}_{ef}$
@@
 21:  end for
-22:
-23:  Sub-thread 1:
-24:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q},\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g}$
-25:  process through the spinal cord module
-26:  update $\bm{\tau}$
-27:
-28:  Sub-thread 2:
-29:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
-30:  process through the brainstem module
-31:  update $\bm{K}_{V},\bm{K}_{P}$
-32:
-33:  Sub-thread 3:
-34:  read $\bm{T}_{{\text{cb}}},\bm{F}_{ef}$
-35:  process through the thalamus module
-36:  update $\bm{\tau}_{g},\bm{w}$
-37:
-38:  Sub-thread 4:
-39:  read $\bm{q}$
-40:  process through the cerebellum module
-41:  update $\bm{T}_{{\text{cb}}}$
-III Module network design
+22:  Sub-thread 1:
+23:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q},\bm{K}_{V},\bm{K}_{P},\bm{\tau}_{g}$
+24:  process through the spinal cord module
+25:  update $\bm{\tau}$
+26:  Sub-thread 2:
+27:  read $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
+28:  process through the brainstem module
+29:  update $\bm{K}_{V},\bm{K}_{P}$
+30:  Sub-thread 3:
+31:  read $\bm{T}_{{\text{cb}}},\bm{F}_{ef}$
+32:  process through the thalamus module
+33:  update $\bm{\tau}_{g}$
+34:  Sub-thread 4:
+35:  read $\bm{q}$
+36:  process through the cerebellum module
+37:  update $\bm{T}_{{\text{cb}}}$
+III Module design
 Building on the framework introduced in the previous section, this chapter focuses on the implementation of each module using SNNs.
 The cerebral cortex module, which serves only for trajectory storage and playback, is therefore excluded from the discussion.
-Since high-performance robotic arms are commonly designed with seven DoF, all implementations in this chapter target a 7-DoF arm.
-For manipulators with different DoF, the same design principles can be applied by appropriately scaling the neural network architecture of each module.
+All implementations in this chapter target a 7-DOF arm, i.e., $n=7$.
 III-A Spiking neuron
-SNN employs spiking neurons as their fundamental computational units.
-A variety of spiking neuron models have been proposed in the literature, such as the Hodgkin–Huxley model [hodgkin1952quantitative] and the Leaky Integrate-and-Fire (LIF) model along with its derivatives [burkitt2006review].
-The LIF model is of particular interest here, since it not only captures the dynamic process of membrane potential accumulation and leakage, but also maintains relatively high computational efficiency [stein1965theoretical].
-Therefore, we will focus on the LIF model.
+SNN employs spiking neurons as its computation units.
+A variety of spiking neuron models have been proposed in the literature, such as the Hodgkin–Huxley model [22] and the Leaky Integrate-and-Fire (LIF) model along with its derivatives [7].
+The LIF model is of particular interest here, since it not only captures the dynamic process of membrane potential accumulation and leakage, but also maintains relatively high computational efficiency [41].
 The input current signal generated by a spike train $S(t)=\sum_{f}s(t-t_{f}),f=1,2,\cdots$ acting on a neuron can be described as
-$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)\exp\left(-s/\tau_{s}\right)\text{d}s}$ (7)
-where $\tau_{s}$ denotes the synaptic time constant. The membrane potential $u$ of the neuron evolves according to the following dynamics:
-$\tau_{m}\frac{\text{d}u}{\text{d}t}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
-where $\tau_{m}=RC$ is the membrane time constant determined by the membrane resistance $R$ and capacitance $C$.
+$i\left(t\right)=\int_{0}^{\infty}{S\left(s-t\right)\exp\left(-s/\tau_{\text{s}}\right)\text{d}s}$ (7)
+where $\tau_{\text{s}}$ denotes the synaptic time constant. The membrane potential $u(t)$ of the neuron evolves according to the following dynamics:
+$\tau_{\text{mb}}\frac{\text{d}u}{\text{d}t}=u_{\text{reset}}-u\left(t\right)+R\left(i_{0}\left(t\right)+\sum w_{j}i_{j}\left(t\right)\right)$ (8)
+where $\tau_{\text{mb}}=RC$ is the membrane time constant determined by the membrane resistance $R$ and capacitance $C$.
 Here, $u_{\text{reset}}$ is the reset potential after firing, $i_{0}(t)$ represents the external current driving the neuron, $i_{j}(t)$ denotes the input current from the $j$-th synapse, and $w_{j}$ is the synaptic weight associated with the $j$-th synapse.
@@
 Therefore, the operational process of a spiking neuron can be interpreted as follows:
-1.When the membrane potential $u(t)$ exceeds the firing threshold $u_{\text{fire}}$, the neuron emits a spike $s(t)$, expressed as
+1. When the membrane potential $u(t)$ exceeds the firing threshold $u_{\text{fire}}$, the neuron emits a spike $s(t)$, expressed as
 $s(t)=\Theta(u(t)-u_{\text{fire}})$ (9)
@@
 When the membrane potential $u(t)$ remains below the threshold $u_{\text{fire}}$, no spike is generated.
-2.At each time step, the differential equation (8) is updated, where the membrane potential decays by $\lvert u_{\text{reset}}-u(t)\rvert$ (representing the forgetting effect over time) and increases by $R\left(i_{0}(t)+\sum w_{j}i_{j}(t)\right)$ (representing the stimulation effect induced by external inputs).
-3.The process then returns to Step 1.
-In our framework, we adopt a backpropagation-based training method[8891809] for SNN.
+2. Update equation (8).
+3. Return to Step 1.
+In our framework, we adopt a back-propagation-based training method[31] for SNN.
 Since the activation function of the neuron in equation (10) is a discontinuous function, directly computing its derivative can lead to network instability.
 Consequently, various surrogate gradient methods have been proposed [31].
-A common approach is to use the original step function $y=\Theta(a)$ during the forward pass, while replacing its derivative with $\theta^{\prime}(a)=\sigma^{\prime}(a)$ instead of $\Theta^{\prime}(a)$ during backpropagation, where $\sigma(a)$ is referred to as the surrogate activation function.
-Typically, $\sigma(a)$ is a smooth and continuous function that resembles the shape of $\Theta(a)$ while enabling stable gradient computation.
+A common approach is to use the original step function $\Theta(x)$ during the forward pass, while replacing its derivative with $\sigma^{\prime}(x)$ instead of $\Theta^{\prime}(x)$ during back propagation, where $\sigma(x)$ is referred to as the surrogate activation function.
+Typically, $\sigma(x)$ is a smooth and continuous function that resembles the shape of $\Theta(x)$ while enabling stable gradient computation.
 As illustrated in Fig.1c, the framework employs three types of neurons across all modules, represented in gray, yellow, and green.
@@
 The yellow neurons represent LIF neurons, which constitute the core units of SNNs and are responsible for information storage, processing, and learning.
-The green neurons correspond to non-spiking LIF neurons, implemented by setting the membrane time constant $\tau_{m}$ of standard LIF neurons to infinity.
+The green neurons correspond to non-spiking LIF neurons, implemented by setting the firing threshold $u_{\text{fire}}$ of standard LIF neurons to infinity.
 Unlike spiking neurons, their output is the continuous membrane potential $u(t)$ rather than discrete spikes, thereby enabling spike-to-continuous signal conversion.
-These neurons serve as the output layer of each module, connecting to the descending pathway, and, in addition to the functional capabilities of the yellow LIF neurons, they also perform information decoding and output.
+These neurons, connecting to the descending pathway, serve as the decoding layer of each module.
 III-B Cerebellum module
@@
 Therefore, compared with unsupervised learning approaches, training a neural network to perform gravity compensation via imitation provides a faster and more efficient solution.
-Under the condition of a known robotic arm dynamics model, the gravity compensation torque represents a nonlinear mapping from joint positions to joint torques.
-In the machine learning domain, the task of fitting a curve to observed data under supervision is referred to as a regression task, for which extensive research and mature methods already exist[henkes2024spiking, robotics13090126].
-Figure 2: (a) Structure of one network unit in the cerebellum module during training. (b) Structure of cerebellum network during prediction.
-As shown in Fig.2a, to implement the cerebellum module, a four-layer recurrent SNN is designed as a network unit of the module, analogous to the “microcomplex” proposed in [abadia2019robot, abadia2021cerebellar].
-The input variable $\bm{q}\in\mathbb{R}^{7}$ corresponds to the actual joint positions of the robotic arm, and the output $\bm{\tau}_{\text{cb}}\in\mathbb{R}^{7}$ represents the gravity compensation torques under a given load condition.
+Under the condition of a known robotic arm dynamic model, the gravity compensation torque represents a nonlinear mapping from joint positions to joint torques.
+In the machine learning domain, the task of fitting a curve to observed data under supervision is referred to as a regression task, for which extensive research and mature methods already exist[21, 42].
+As shown in Fig.2a, to implement the cerebellum module, a four-layer recurrent SNN inspired by [2, 3] is designed to store one motor pattern corresponding to a particular load mass.
+The input variable $\bm{q}\in\mathbb{R}^{7}$ corresponds to the actual joint positions of the robotic arm, and the output $\bm{\tau}_{\text{cb}}\in\mathbb{R}^{7}$ represents the gravity compensation torque under a given load condition.
+Figure 2: Structure of the cerebellum module. (a) In training. (b) In prediction.
 Variables from the ascending pathway first arrive at the mossy fiber (MF) layer, which is fully connected to the granule cell (GC) layer consisting of 300 LIF neurons.
@@
 The output of the PC layer is fully connected to the deep cerebellar nuclei (DCN) layer, which consists of 140 non-spiking LIF neurons.
-The decoding of the DCN layer is inspired by population coding observed in the motor cortex of animals [georgopoulos1986neuronal].
-Studies have shown that during arm movement in primates, the activity of any single motor cortex neuron does not directly indicate the actual movement direction; however, a weighted sum of vectors corresponding to the preferred directions of a population of neurons yields a vector aligned with the movement direction.
+The decoding of the DCN layer is inspired by population coding in the animal motor cortex [16].
+During primate arm movements, individual motor cortex neurons do not directly represent movement direction; rather, the movement direction emerges from a weighted vector sum over the preferred directions of a neuronal population.
 Analogously, the 140 DCN neurons are grouped in sets of 20, and their membrane potentials are averaged to obtain a 7-dimensional torque output $\bm{\tau}_{\text{cb}}$, i.e.,
-$\tau_{\text{cb},i}=\sum_{j=20\cdot(i-1)}^{20\cdot i}u_{\text{DCN},j}$ (11)
+$\tau_{\text{cb},i}=\sum_{j=20\cdot(i-1)+1}^{20\cdot i}u_{\text{DCN},j}$ (11)
 where $\tau_{\text{cb},i},i=1,\cdots,7$ represents the torque for each joint and $u_{\text{DCN},j},j=1,\cdots,140$ denotes the membrane potential of each DCN neuron.
-The network employs rate-based encoding[gerstner2002spiking], with a time window of length $T$.
-The training data $\bm{\tau}_{\text{ref}}$ are generated using the robotic arm’s dynamics model.
+The network employs rate-based encoding[17], with a time window of length $T$.
+The training data $\bm{\tau}_{\text{ref}}$ are generated using the robotic arm’s dynamic model.
 Under the condition that the training data adequately cover the workspace, this regression-based approach is sufficient to achieve effective gravity compensation.
-The network is trained in a supervised manner, with the objective of minimizing the error between the torques output by the network and the reference torques computed from the dynamics model.
+The network is trained in a supervised manner, with the objective of minimizing the error between the network output and the reference torque.
 Structurally, this is reflected in the DCN layer neurons receiving both excitatory inputs from the PC layer neurons and inhibitory inputs from the climbing fiber (CF) neurons.
@@
 $\text{MSE}=\frac{\sum_{i=1}^{7}(\tau_{\text{cb,i}}-\tau_{\text{ref,i}})^{2}}{7}$ (12)
-The weights are then updated via backpropagation, thereby completing the training process; detailed procedures are provided in Algorithm 2.
-Algorithm 2 Training process of one network unit in the cerebellum module
+The weights are then updated via back propagation, thereby completing the training process.
+Detailed training procedure is provided in Algorithm 2.
+After training, the network can operate without reference torques, relying solely on the learned synaptic weights to generate gravity compensation torque.
+Algorithm 2 Training process of network for one motor pattern
 1:  initialize load $M$, trajectory length $L$, batch size bsbs, learning rate lrlr, momentum momo, epochs epep, random noise $\bm{n}$, network loss lossloss
@@
 TABLE I: Training and neuron parameters of the cerebellum module
-Training parameter Value Neuron parameter Value
+Parameter Value Parameter Value
 Epochs 50 each trajectory $u_{\text{reset}}$ GC: 0.0
@@
 Optimizer SGD DCN: $\infty$
-$\theta_{0}$ $\frac{\pi}{6}$ $\tau_{m}$ GC: 10.0
+$\theta_{0}$ $\frac{\pi}{6}$ $\tau_{\text{mb}}$ GC: 10.0
 $R_{c}$ 0.14 PC: 10.0
@@
 $(x_{1},y_{1},z_{1})$ (0.63,-0.11,0.3) $T$ 10
-After training, the network can operate without reference torques, with CF neurons no longer required, relying solely on the learned synaptic weights to generate gravity compensation torques.
-To accommodate different loads, two network units are trained separately under light (no load) and heavy (3 kg load) modes using the supervised learning procedure described in Algorithm 2.
-These trained networks are then combined to form the complete cerebellum module structure illustrated in Fig.
+To realize motor patterns for different load masses, two networks, as depicted in Fig.
+2a, are trained separately in light pattern (with no load) and in heavy pattern (with 3 kg load) using the supervised learning procedure described in Algorithm 2.
+These trained networks are then combined to form the complete cerebellum module, as illustrated in Fig.
 2b.
 In summary, after training, the cerebellum module receives as input the actual joint positions of the robotic arm, $\bm{q}\in\mathbb{R}^{7}$, and outputs a gravity compensation matrix composed of two torque vectors, i.e., $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$, $\bm{\tau}_{\text{cb,L}}\in\mathbb{R}^{7}$ and $\bm{\tau}_{\text{cb,H}}\in\mathbb{R}^{7}$.
-This matrix corresponds to the learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy conditions.
-The outputs are subsequently integrated by the thalamus module to achieve gravity compensation for arbitrary loads within a specified range.
-During trajectory tracking, the prediction process of the cerebellum module is described in Algorithm 3, which can be implemented as Sub-thread 4 in Algorithm 1.
+This matrix corresponds to the two learned patterns of the cerebellum module and provides gravity compensation for trajectory tracking tasks under both light and heavy pattern.
+During trajectory tracking tasks, the prediction process of the cerebellum module is described in Algorithm 3, which can be implemented as Sub-thread 4 in Algorithm 1.
 Algorithm 3 Prediction process of the cerebellum module
-1:  load the network for light and heavy load conditions
+1:  load the cerebellum module network
 2:  initialize gravity compensation matrix $\bm{T}_{\text{cb}}$
 3:  read joint position $\bm{q}$
-4:  feed $\bm{q}$ into the two trained network unit, and generate two output $\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}}$
+4:  feed $\bm{q}$ into the cerebellum module network, and generate two output $\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}}$
 5:  update $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})$
 III-C Thalamus module
-As described in the previous sub-section, although we trained two neural network units under different end-effector load conditions to form the cerebellum module—thereby generating gravity compensation torques corresponding to two specific load cases—the cerebellum module alone remains insufficient for handling arbitrary load.
+As described in the previous sub-section, although we trained two neural networks to form the cerebellum module—thereby generating gravity compensation torques corresponding to two specific load masses—the cerebellum module alone remains insufficient for handling arbitrary load.
 To address this limitation, we developed the thalamus module.
-By utilizing torque sensor data at the robotic arm’s end-effector, the thalamus module network is able to “sense” the weight of the object, akin to human perception, and accordingly assign adaptive weights to the two gravity compensation torques produced by the cerebellum module’s motor patterns via reinforcement learning.
-This mechanism enables effective compensation for loads of varying mass and shape.
-Figure 3: Structure of thalamus module.
+By utilizing torque sensor data from the robotic arm’s end-effector, the thalamus module network is able to “sense” the mass of the object, akin to human perception, and accordingly assign adaptive weights to the two gravity compensation torques produced by the cerebellum module’s motor patterns via reinforcement learning.
 The thalamus module contains five layers of neurons, as illustrated in Fig.
 3.
-Its input $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$ and $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\in\mathbb{R}^{7\times 2}$, corresponds to the six-dimensional force/torque sensor data measured at the end-effector, and the cerebellum module output, respectively.
-Its output, $\bm{w}=(w_{1},w_{2})^{T}\in\mathbb{R}^{2}$, represents the weight coefficients assigned to the two gravity compensation torques provided by the cerebellum module.
+Its input $\bm{F}_{\text{ef}}\in\mathbb{R}^{6}$ and $\bm{T}_{\text{cb}}=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\in\mathbb{R}^{7\times 2}$, correspond to the six-dimensional force/torque sensor data measured at the end-effector, and the cerebellum module output, respectively.
+Its output, $\bm{\tau}_{g}\in\mathbb{R}^{7}$, represents the gravity compensation torque for arbitrary load mass within a certain range.
+Figure 3: Structure of the thalamus module.
 The input layer is fully connected to a hidden layer consisting of five LIF neurons, which in turn connects to an output layer of two non-spiking LIF neurons.
-The membrane potentials of the output neurons are then passed through a Softmax function to produce the two weight coefficients.
-The loss function designed for this network comprises two terms, minimizing both the joint position error and the joint velocity error, i.e.,
+The membrane potential of the output neurons are then passed through a Softmax function to produce the two weight coefficients.
+The network is trained via reinforcement learning, and the loss function designed for this network comprises two terms, minimizing both the joint position error and the joint velocity error, i.e.,
 $\text{Loss}=\sum_{i=1}^{7}100*(q_{d,i}-q_{i})+\sum_{i=1}^{7}10^{-4}\cdot(\dot{q}_{d,i}-\dot{q}_{i})$ (16)
@@
 Momentum 0.5 Filter layer: 0.0
-$\tau_{m}$ Hidden layer: 5.0 $u_{\text{fire}}$ Hidden layer: 0.01
+$\tau_{\text{mb}}$ Hidden layer: 5.0 $u_{\text{fire}}$ Hidden layer: 0.01
 Output layer: 5.0 Output layer: $\infty$
 Filter layer: 5.0 Filter layer: $\infty$
-After obtaining $\bm{w}$ from the output layer, it is then combined with data from the connecting fiber (linking the thalamus module and the cerebellum module) to generate a weighted output via matrix multiplication, yielding
-$\begin{split}\delta\bm{u}_{\text{ft}}(t)&=(\delta u_{\text{ft,1}}(t),\delta u_{\text{ft,2}}(t),\cdots,\delta u_{\text{ft,7}}(t))^{T}\\ &=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\cdot\bm{w}=w_{1}\bm{\tau}_{\text{cb,L}}+w_{2}\bm{\tau}_{\text{cb,H}}\end{split}$ (17)
-This weighted result is not directly used as the output; instead, it is injected as an increment to the membrane potentials of seven non-spiking LIF neurons that serve as a filter layer.
+After obtaining $\bm{w}$ from the output layer, it is then combined with data from the connecting fiber (linking the thalamus module and the cerebellum module) to generate a weighted output, yielding
+$\begin{split}\delta\bm{u}_{\text{ft}}(t)&=(\delta u_{\text{ft,1}}(t),\delta u_{\text{ft,2}}(t),\cdots,\delta u_{\text{ft,7}}(t))^{\top}\\ &=(\bm{\tau}_{\text{cb,L}},\bm{\tau}_{\text{cb,H}})\cdot\bm{w}=w_{1}\bm{\tau}_{\text{cb,L}}+w_{2}\bm{\tau}_{\text{cb,H}}\end{split}$ (17)
+This would then be injected as an increment to the membrane potentials of seven non-spiking LIF neurons that serve as a filter layer.
 Specifically, the membrane potential of each filtering-layer neuron at time $t$ is given by
-$u_{\text{ft},i}(t)$ $=u_{\text{ft},i}(t-1)-\frac{u_{\text{ft},i}(t-1)-u_{\text{reset}}}{\tau_{m}}$ (18)
-$+\frac{\delta u_{\text{ft},i}(t)}{\tau_{m}},i=1,\cdots,7$
-Finally we take the membrane potential of each filter-layer neuron at time $t$ as the gravity compensation torque, i.e., $\bm{\tau}_{g}=\bm{u}_{\text{ft}}(t)=(u_{\text{ft},1}(t),u_{\text{ft},2}(t),\cdots,u_{\text{ft},7}(t))$, which is then transmitted to the spinal cord module for further processing.
+$u_{\text{ft},i}(t)$ $=u_{\text{ft},i}(t-1)-\frac{u_{\text{ft},i}(t-1)-u_{\text{reset}}}{\tau_{\text{mb}}}$ (18)
+$+\frac{\delta u_{\text{ft},i}(t)}{\tau_{\text{mb}}},i=1,\cdots,7$
+Finally we take the membrane potential of each filter-layer neuron at time $t$ as the gravity compensation torque, i.e., $\bm{\tau}_{g}=\bm{u}_{\text{ft}}(t)=(u_{\text{ft},1}(t),u_{\text{ft},2}(t),\cdots,u_{\text{ft},7}(t))^{\top}$.
 The parameters of the neurons in each layer of this network, as well as the training parameters, are summarized in Table II.
@@
 Algorithm 4 Training process of the thalamus module
-1:  initialize gravity compensation torque $\bm{\tau}_{g}$, weight coefficients $\bm{w}$, network loss lossloss, optimizer $optim(lr,mo)$
+1:  initialize gravity compensation torque $\bm{\tau}_{g}$, weight coefficients $\bm{w}$, learning rate lrlr, momentum momo, network loss lossloss, optimizer $optim(lr,mo)$
 2:  read $\bm{T}_{{\text{cb}}},\bm{F}_{\text{ef}}$
@@
 III-D Brainstem module
-As illustrated in Fig.
-1c, the primary role of the brainstem module is to dynamically select appropriate feedback control coefficients during operation through learning, thereby compensating for the nonlinear dynamics of the robotic arm and the coupling effects among joints.
-The network is implemented as a three-layer fully connected SNN.
-Its input variable $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ is a concatenated vector consisting of the desired joint positions and velocities, as well as the actual joint positions and velocities.
-The output $\bm{K}=(k_{p,1},k_{p,2},\cdots,k_{p,7},k_{v,1},k_{v,2},\cdots,k_{v,7})\in\mathbb{R}^{14}$ corresponds to the proportional and derivative coefficients for each of the seven joints.
+As stated in section II-A4, the brainstem module is to provide appropriate feedback control coefficients during operation to compensate for the nonlinear dynamics of the robotic arm and the coupling effects among joints.
+Its network is implemented as a three-layer fully connected SNN, as illustrated in Fig.
+1c.
+The input variable $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ is a concatenated vector consisting of the desired joint positions and velocities, as well as the actual joint positions and velocities.
+The output $\bm{K}_{P}=diag(k_{p,1},k_{p,2},\cdots,k_{p,7})$, $\bm{K}_{V}=diag(k_{v,1},k_{v,2},\cdots,k_{v,7})$ corresponds to the proportional and derivative coefficients for each of the seven joints.
 The input vector is fully connected to a hidden layer of 10 LIF neurons, which in turn is fully connected to an output layer of 14 non-spiking LIF neurons.
@@
 […]
-In this formulation, the first term minimizes the position error and the second term minimizes the velocity error, ensuring accurate trajectory tracking, while the third term penalizes torque usage to reduce joint energy consumption and prevent abrupt fluctuations in joint torque.
-Through this design, the network is able to learn feedback control coefficients that balance trajectory-tracking accuracy with reduced joint wear and energy cost.
-Similar to the cerebellum module, the brainstem module adopts rate–based encoding with a time window length of $T$, and the approximate gradient function is chosen as the sigmoid function.
+In this formulation, the first term and second term minimize position and tracking error to ensure accurate trajectory tracking, while the third term penalizes torque usage to reduce joint energy consumption.
+Similar to the cerebellum module, the brainstem module adopts rate–based encoding with a time window length of $T$, and the surrogate gradient function is chosen as the sigmoid function.
 After decoding through the non-spiking LIF neurons in the output layer, we obtain a set of membrane potentials $\bm{u}_{\text{bs}}=(u_{\text{bs},1},\cdots,u_{\text{bs},14})\in\mathbb{R}^{14}$.
@@
 $\nu$ 5.0 Output layer: $\infty$
-(a,b) (20,2) $\tau_{m}$ Hidden layer: 5.0
+(a,b) (20,2) $\tau_{\text{mb}}$ Hidden layer: 5.0
 $T$ 15 Output layer: 5.0
@@
 Algorithm 5 Training process of the brainstem module
-1:  initialize coefficients $\bm{K}_{V},\bm{K}_{P}$, network loss lossloss, optimizer $optim(lr,mo)$
+1:  initialize coefficients $\bm{K}_{V},\bm{K}_{P}$, learning rate lrlr, momentum momo, network loss lossloss, optimizer $optim(lr,mo)$
 2:  read joint position and velocity data $\bm{\dot{q}}_{d},\bm{q}_{d},\bm{\dot{q}},\bm{q}$
@@
 III-E Spinal cord module
-The primary function of the spinal cord module is to integrate proprioceptive information with the outputs of the aforementioned higher-level motor control modules, thereby generating real-time motion control commands.
-Its structural design is illustrated in Fig.
+According to section II-A5, the spinal cord module is designed to integrate proprioceptive information and the outputs of the aforementioned modules to generate real-time motor control commands.
+For high-performance robotic arms, torque control typically requires the control algorithm to achieve frequencies of 1 kHz or even higher.
+Consequently, the spinal cord module must exhibit very high computational efficiency.
+The structure of this module is illustrated in Fig.
 1c.
-For high-performance robotic manipulators, torque control typically requires the control algorithm to achieve frequencies of 1000 Hz, 1200 Hz, or even higher.
-Consequently, the spinal cord module must exhibit very high computational efficiency.
-In our design, the input variables $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ consist of the planned positions and velocities as well as the actual positions and velocities of all joints, while the output $\bm{\tau}\in\mathbb{R}^{7}$ represents the control torque for each joint.
+The input variables $\bm{q}_{d},\bm{\dot{q}}_{d},\bm{q},\bm{\dot{q}}]\in\mathbb{R}^{28}$ consist of the planned positions and velocities as well as the actual positions and velocities of seven joints, while the output $\bm{\tau}\in\mathbb{R}^{7}$ represents the control torque for each joint.
 After entering the spinal cord module, the planned trajectory and proprioceptive feedback are first encoded into discrete spike trains, with each joint being processed independently.
-Specifically, the planned position and velocity signals $q_{d,i},\dot{q}_{d,i},i=1,\cdots,7$ serve to excite the input fibers, whereas the actual proprioceptive feedback $q_{i},\dot{q}_{i},i=1,\cdots,7$ inhibits the input fibers.
+Specifically, the planned position and velocity signals $q_{d,i},\dot{q}_{d,i},i=1,\cdots,7$ excite the input fibers, whereas the actual proprioceptive feedback $q_{i},\dot{q}_{i},i=1,\cdots,7$ inhibit the input fibers.
 Mathematically, this process is equivalent to computing their difference.
 The outputs of the input fibers are subsequently passed into the Poisson encoding stage, given by
-$s_{q,i,j}=\begin{cases}0,&-u_{j}<\frac{q_{d,i}-q_{i}}{\Delta q_{i}}<u_{j},j=1,\cdots,n\\ 1,&\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=1,\cdots,\frac{n}{2}\\ -1,&-\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=\frac{n}{2},\cdots,n\end{cases}$ (22)
-$s_{\dot{q},i,j}=\begin{cases}0,&-u_{j}<\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}<u_{j},j=1,\cdots,n\\ 1,&\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=1,\cdots,\frac{n}{2}\\ -1,&-\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=\frac{n}{2},\cdots,n\end{cases}$ (23)
-Here, $u_{j}\sim U(0,1),j=1,\cdots,n$ denotes $n$ independent uniformly distributed random variables, ${s_{q,i,j}},i=1,\cdots,7,j=1,\cdots,n$ represents the spike trains encoding the position information of each joint at the current time step, and ${s_{\dot{q},i,j}},i=1,\cdots,7,j=1,\cdots,n$ denotes the spike trains encoding the velocity information.
-The terms $\Delta q_{i}$ and $\Delta\dot{q}_{i},i=1,\cdots,7$ correspond to the maximum allowable position and velocity errors for each joint, respectively.
-From an intuitive perspective, each input fiber is connected to a group of $n$ submodules, where the first half $\frac{n}{2}$ are responsible for processing positive inputs and the second half $\frac{n}{2}$ handle negative inputs.
+$s_{q,i,j}=\begin{cases}0,&-u_{j}<\frac{q_{d,i}-q_{i}}{\Delta q_{i}}<u_{j},j=1,\cdots,k\\ 1,&\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=1,\cdots,\frac{k}{2}\\ -1,&-\frac{q_{d,i}-q_{i}}{\Delta q_{i}}\geq u_{j},j=\frac{k}{2},\cdots,k\end{cases}$ (22)
+$s_{\dot{q},i,j}=\begin{cases}0,&-u_{j}<\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}<u_{j},j=1,\cdots,k\\ 1,&\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=1,\cdots,\frac{k}{2}\\ -1,&-\frac{\dot{q}_{d,i}-\dot{q}_{i}}{\Delta\dot{q}_{i}}\geq u_{j},j=\frac{k}{2},\cdots,k\end{cases}$ (23)
+Here, $u_{j}\sim U(0,1),j=1,\cdots,k$ denotes $k$ independent uniformly distributed random variables, ${s_{q,i,j}},{s_{\dot{q},i,j}},i=1,\cdots,7,j=1,\cdots,k$ denote the spike trains encoding the position and velocity information of each joint at each time step, respectively.
+The terms $\Delta q_{i}$ and $\Delta\dot{q}_{i},i=1,\cdots,7$ correspond to the maximum position and velocity errors for each joint, respectively.
+From an intuitive perspective, each input fiber is connected to a group of $k$ neurons, where the first half $\frac{k}{2}$ are responsible for processing positive inputs and the second half $\frac{k}{2}$ handle negative inputs.
 This structure corresponds to the agonist–antagonist pairing characteristic of human muscles.
-Each submodule generates a uniformly distributed random number in the range $[0,1]$.
-When the output of the input fiber is positive, the first $\frac{n}{2}$ neurons are encoded sequentially: if the absolute value of the input exceeds the random number, the encoding result is set to $1$; otherwise, it is 0.
-Conversely, when the output of the input fiber is negative, the second $\frac{n}{2}$ neurons are encoded sequentially: if the absolute value of the input exceeds the random number, the encoding result is set to 1-1; otherwise, it is 0.
-This encoding scheme can be implemented in hardware using simple logic gate circuits, thereby achieving very high computational efficiency.
-The encoding precision is influenced by three parameters: $\Delta q_{i}$, $\Delta\dot{q}_{i}$, $i=1,\cdots,7$, and $n$.
-Smaller values of $\Delta q_{i}$ and $\Delta\dot{q}_{i}$ lead to higher encoding precision, but reduce the range of representable errors; a larger $n$ increases encoding precision but reduces computational efficiency.
-Following this encoding process, two sets of discrete spike trains are generated—one representing position information and the other representing velocity information.
-These spike trains are then transmitted to stimulate the activity of non-spiking LIF neurons in the output layer, with stimulation strength modulated by the feedback control gains $k_{p,i}$ and $k_{v,i}$, $i=1,\cdots,7$ provided by the brainstem module, as illustrated by the blue and red dashed lines, respectively.
-Accordingly, at each time step, the membrane potential increments of the output non-spiking LIF neurons can be expressed as
+Each neuron generates a uniformly distributed random number in the range of $[0,1]$.
+When the output of the input fiber is positive, the first $\frac{k}{2}$ neurons are activated: if the absolute value of each neuron’s input exceeds the random number, the encoding result is set to $1$; otherwise, it is 0.
+Conversely, when the output of the input fiber is negative, the second $\frac{k}{2}$ neurons are activated: if the absolute value of each neuron’s input exceeds the random number, the encoding result is set to 1-1; otherwise, it is 0.
+The encoding effect is influenced by three parameters: $\Delta q_{i}$, $\Delta\dot{q}_{i}$, $i=1,\cdots,7$, and $k$.
+Smaller values of $\Delta q_{i}$ and $\Delta\dot{q}_{i}$ lead to higher encoding precision, but reduce the dynamic range; a larger $k$ increases encoding precision but reduces computational efficiency.
+Following this encoding process, two sets of spike trains are generated—one representing position information and the other representing velocity information.
+These spike trains are then transmitted to stimulate the activity of the non-spiking LIF neuron for each joint, with stimulation strength modulated by the feedback control coefficients $k_{p,i}$ and $k_{v,i}$, $i=1,\cdots,7$ provided by the brainstem module.
+Accordingly, at each time step, the membrane potential increment of the output non-spiking LIF neurons can be expressed as
 $\delta u_{\text{sp,i}}(t)$ $=\tau_{g,i}+k_{p,i}\cdot\sum_{j=1}^{100}s_{q,i,j}$ (24)
 $+k_{v,i}\cdot\sum_{j=1}^{100}s_{\dot{q},i,j},i=1,\cdots,7$
-Here, $\bm{\tau}_{g}=(\tau_{g,1},\tau_{g,2},\cdots,\tau_{g,7})$ denotes the gravity compensation torque generated by the thalamus module.
+Here, $\bm{\tau}_{g}=(\tau_{g,1},\tau_{g,2},\cdots,\tau_{g,7})^{\top}$ denotes the gravity compensation torque generated by the thalamus module.
 Consequently, the membrane potentials at each time step can be written as
-$u_{\text{sp},i}(t)$ $=u_{\text{sp},i}(t-1)-\frac{u_{\text{sp},i}(t-1)-u_{\text{reset}}}{\tau_{m}}$ (25)
-$+\frac{\delta u_{\text{sp},i}(t)}{\tau_{m}},i=1,\cdots,7$
-Finally, similar to the thalamus module, we take $\bm{\tau}=\bm{u}_{\text{sp}}=(u_{\text{sp},1}(t),u_{\text{sp},2}(t),\cdots,u_{\text{sp},7}(t))$ as the final control torque, which will be sent in real time to the actuators.
-The neuronal parameters of the spinal cord module are summarized in Table IV.
+$u_{\text{sp},i}(t)$ $=u_{\text{sp},i}(t-1)-\frac{u_{\text{sp},i}(t-1)-u_{\text{reset}}}{\tau_{\text{mb}}}$ (25)
+$+\frac{\delta u_{\text{sp},i}(t)}{\tau_{\text{mb}}},i=1,\cdots,7$
+Finally, similar to the thalamus module, we take $\bm{\tau}=\bm{u}_{\text{sp}}=(u_{\text{sp},1}(t),u_{\text{sp},2}(t),\cdots,u_{\text{sp},7}(t))^{\top}$ as the control torque to be sent in real time to the actuators.
+The neuronal and encoding parameters of the spinal cord module are summarized in Table IV.
 TABLE IV: Neuron parameters of the spinal cord module
@@
 $u_{\text{fire}}$ $\infty$ $\Delta\dot{q}_{i},i=1,\cdots,7$ 0.5
-$\tau_{m}$ 2.0 n 100
-Through this mechanism, the output torque at each time step incorporates both the current information and the historical information over a preceding time window, thereby making it more suitable than ANN for time-continuous tasks such as robotic control.
+$\tau_{\text{mb}}$ 2.0 k 100
 During trajectory tracking tasks, the operational process of the spinal cord module is illustrated in Algorithm 6, which can be implemented as Sub-thread 1 within Algorithm1.
@@
 7:  read potential and update $\bm{\tau}$
-IV Simulation analysis
-To evaluate the rationality and effectiveness of the proposed framework, we conduct three simulation experiments that demonstrate contributions (1)–(3).
-In section IV-A and IV-B, we validate the design of the modules and in section IV-C, the design of the framework.
-The algorithm is implemented in Python (version 3.12.4), and all neural networks are constructed using PyTorch [paszke2019pytorch] (version 2.4.0).
-The spiking neurons involved in the algorithm are implemented with the SNN learning framework SpikingJelly [doi:10.1126/sciadv.adi1480] (version 0.0.0.0.14).
-The simulation environment employed is PyBullet [coumans2021] (version 3.2.6), an open-source robotics simulation tool whose underlying physics engine is based on the open-source Bullet Physics SDK.
-The controlled robotic arm consists of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector.
-The control cycle of each layer is set as $T_{1}=1$ms, $T_{2}=10$ms, $T_{3}=50$ms.
+IV Simulation results
+In this section, we conduct simulation experiments to validate that the module designs proposed in Section III satisfy the requirements outlined in Section II, and to evaluate the resilience of the proposed framework.
+The framework is implemented in Python (version 3.12.4), and all modules are constructed using PyTorch [34] (version 2.4.0).
+The spiking neurons involved in the framework are implemented with the SNN learning framework SpikingJelly [14] (version 0.0.0.0.14).
+The simulation environment employed is PyBullet [11] (version 3.2.6), an open-source robotics simulation tool whose underlying physics engine is based on the open-source Bullet Physics SDK.
+The controlled robotic arm is a Flexiv Rizon 4s consisting of seven actuated joints, seven links and a six-dimensional torque sensor mounted at the flange of the end-effector.
+The framework runs on a multi-core CPU, and the control cycle of the first-order, second-order and third-order level is set as $T_{1}=1$ms, $T_{2}=10$ms, $T_{3}=50$ms, respectively.
 The simulation time step is set to 1 ms, and the control frequency is configured to 1 kHz.
-In all simulation experiments presented in this section, trajectory tracking accuracy is evaluated using the root-mean-square error (RMSE) of the joint positions.
+In all simulation experiments presented in this section, trajectory tracking accuracy is evaluated using the root mean square error (RMSE) of the joint positions.
 The RMSE at each time step is defined as
@@
 while the average RMSE over an entire trajectory is defined as
-$\text{RMSE}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}(t)$ (27)
+$\text{RMSE}=\frac{1}{L}\sum_{t=1}^{L}\text{RMSE}(t)$ (27)
+where $L$ is the trajectory length.
 Three types of trajectories are employed in the simulations: horizontal circle, inclined circle, and spatial figure-eight.
 The first two trajectories have already been defined in section III-B, while the third trajectory is defined as
-spatial figure-eight: $\begin{cases}x=x+0.5R\sin(\frac{4\pi t}{T})\\ y=y+R\cos(\frac{2\pi t}{T})\\ z=z+z\sin(\frac{2\pi t}{T})\end{cases}$ (28)
-where $x=0.61,y=0,z=0.3,R=0.14,T=3,z=0.08$.
-Figure 4: Average feedback control coefficients of each joint across 3 trajectory tracking tasks.
-IV-A Spinal cord and brainstem module validation
-Physiological studies have demonstrated that joint stiffness and damping in the human arm exhibit an inverse correlation [bennett1992time].
-Specifically, stiffness decreases when a joint begins to move and increases when at rest, while damping decreases as the joint approaches a target and increases at the onset of motion.
-In the proposed framework, the feedback control coefficients $\bm{K}_{V}$ and $\bm{K}_{P}$ generated by the brainstem module correspond precisely to damping and stiffness in feedback control.
+spatial figure-eight: $\begin{cases}x=x_{0}+0.5R_{e}\sin(\frac{4\pi t}{T_{e}})\\ y=y_{0}+R_{e}\cos(\frac{2\pi t}{T_{e}})\\ z=z_{0}+0.08\sin(\frac{2\pi t}{T_{e}})\end{cases}$ (28)
+where $x_{0}=0.61,y_{0}=0,z_{0}=0.3,R_{e}=0.14,T_{e}=3$.
+IV-A Module design validation
 Fig.
-4 illustrates the variations of average $k_{p}$ (blue solid lines) and $k_{v}$ (red dashed lines) for each joint across three trajectories.
-It can be observed that the two parameters generally exhibit opposite trends, as exemplified by joints 3–6 during 3–5 s and joints 4 and 6 during 1–2 s.
-This finding indicates that the brainstem module, through reinforcement learning, adjusts joint stiffness and damping in the spinal module in a manner that exhibits biomimetic characteristics, consistent with our design objectives.
-We also note that this feature is less pronounced in joints closer to the robot’s base, such as joints 1 and 2, likely due to stronger coupling effects from adjacent joints.
-Figure 5: (a) Average tracking error using single mode and two modes across three trajectories. (b) Average weight generated by the thalamus module across three trajectories.
-IV-B Thalamus and cerebellum module validation
+4 illustrates the performance of the cerebellum and thalamus modules.
+As shown in Fig.
+4a, when only the light motor pattern is used, the tracking error increases with increasing load mass, whereas when only the heavy motor pattern is used, the tracking error decreases as the load mass increases.
+This indicates that the cerebellum module effectively encodes two distinct motor patterns: the light pattern corresponds to a load of approximately 1 kg, while the heavy pattern corresponds to a load of approximately 3.5 kg.
+However, neither pattern alone can accommodate arbitrary load masses within the range of 1 kg to 3.5 kg.
+Owing to its online learning capability, the thalamus module can adaptively adjust the weighting coefficients assigned to these two motor patterns.
+As illustrated in Fig.
+4b, the thalamus module assigns a smaller weight to the light pattern and a larger weight to the heavy pattern as the load mass increases.
+Consequently, the weighted combination of the two patterns achieves consistently good performance across the entire load range from 1 kg to 3.5 kg.
+This effect is evident in Fig.
+4a, where the black curve remains below the red and blue curves.
+This adaptive mechanism mirrors human motor control, in which movements are generated through linear combinations of motor primitives [38], thereby highlighting the biomimetic nature of the cerebellum and thalamus modules.
+Overall, these results demonstrate that the designs of both modules satisfy the proposed requirements.
+Figure 4: Performance of the cerebellum and thalamus modules.
+(a) Average RMSE across three trajectories using single motor pattern and two motor patterns.
+(b) Average motor pattern weight coefficients across three trajectories generated by the thalamus module.
+Figure 5: Performance of the brainstem module.
 Fig.
-5a shows that single gravity compensation modes from the cerebellum module are only effective within narrow load ranges (e.g., the light mode around 1 kg and the heavy mode around 3.5 kg), with errors otherwise increasing nearly linearly with load variation.
-In contrast, the thalamus module adaptively synthesizes both modes, maintaining stable joint error across the entire 1–3.5 kg range, thereby validating its design.
+5 illustrates the evolution of the average stiffness coefficients $k_{p,i},i=1,\cdots,7$ and damping coefficients $k_{v,i},i=1,\cdots,7$ for each joint across three trajectories.
+The results show that the brainstem module can adaptively generate feedback control gains for the spinal cord module.
+Notably, the stiffness and damping coefficients of joints 4 and 6 generally exhibit opposite trends.
+Quantitative analysis confirms this observation: the correlation coefficients between $k_{p,4}$ and $k_{v,4}$, and between $k_{p,6}$ and $k_{v,6}$, are 0.5640-0.5640 and 0.8515-0.8515, respectively.
+This behavior is consistent with human motor control.
+Physiological studies have shown that joint stiffness and damping in the human arm are inversely correlated [5], indicating that the brainstem module exhibits biomimetic characteristics.
+Overall, these results demonstrate that the design of the brainstem module satisfies the proposed requirements.
+IV-B Resilience performance evaluation
+Figure 6: Resilience performance under unstructured load changes.
+(a) Evolution of average RMSE across the three trajectories.
+(b) Expanded view of 0-1.5 s.
+(c) Expanded view of 12-14 s.
+(d) Expanded view of 24-25.5 s.
+(e) Expanded view of 0-1.5 s without the brainstem module.
+(f) Expanded view of 0-1.5 s without the thalamus module.
+(g) Expanded view of 0-1.5 s without the cerebellum module.
+Fig.
+6 illustrates the resilience of the proposed framework under unstructured load variations.
 As shown in Fig.
-5b, reinforcement learning enables the thalamus module to adjust mode weights according to load mass, assigning a smaller weight to the light mode and a greater weight to the heavy mode as the load increases.
-This adaptive mechanism mirrors human motor control, where movements are achieved through linear combinations of motor primitives[sanger2000human], thus highlighting the biomimetic nature of the thalamus and cerebellum module.
-Figure 6: Average joint error across the three trajectories under varying load conditions during operation.
-(a) 0-1.5s.
-(b) 12-14s.
-(c) 24-25.5s.
-(d) 0-1.5s without the brainstem module.
-(e) 0-1.5s without the thalamus module.
-(f) 0-1.5s without the cerebellum module.
-IV-C Framework validation
-Fig.
-6 illustrates the average trajectory tracking error across three trajectories under different load conditions.
+6a, when a 1.5 kg load is applied at $t=0$ s, the framework recovers after a learning period of approximately 1.5 s.
+After 12 s of operation, an additional 2 kg load is introduced, and the system again recovers within about 2 s.
+When the added load is removed at 24 s, recovery is achieved within approximately 1.5 s.
+In contrast, cerebellum-inspired networks [2, 3, 1] require nearly 100 s to adapt to a 500 g load change, while CBMC-V2 [32] requires about 15 s to converge after initialization.
+Furthermore, the resilience performance improves over time.
+Although the load change at 12 s is larger than that at 0 s, the resulting tracking error fluctuation during recovery is considerably smaller, as shown in Fig.
+6b and Fig.
+6c.
+Similarly, as illustrated in Fig.
+6d, the fluctuation induced by load removal at 24 s is even smaller, despite the mass change being identical to that at 12 s.
+In contrast, the error fluctuations caused by load variations in [2, 3, 1] do not diminish as operation time increases.
+It can also be observed from Fig.
+6b, Fig.
+6c, and Fig.
+6d that the recovery process following each load change consistently exhibits three distinct error peaks.
+These peaks indicate that the framework converges in a hierarchical manner.
+Specifically, the spinal cord module (first-level) operates at the real-time control frequency and corresponds to the response reflected at each data point of the curve; the brainstem and thalamus modules (second-level) account for the first two peaks; and the cerebellum module (third-level) is responsible for the final peak.
+This stepwise activation pattern is analogous to human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [15].
+To verify that these peaks are associated with the framework’s hierarchical structure, ablation experiments are conducted by selectively disabling individual modules.
 As shown in Fig.
-6a, when a 1.5kg load is applied at 0s, the algorithm converges to a stable state within approximately 1.5s; after 12s of operation, adding an additional 2kg load again leads to convergence within about 2s; and upon removing the added load at 24s, convergence is also achieved within 2s.
-In contrast, the cerebellum-inspired network proposed in [abadia2019robot, abadia2021cerebellar] requires nearly 100s to adapt to a 500g change in load, highlighting the superior dynamic performance and convergence speed of our framework.
-Furthermore, we observe that, compared with cold-start initialization, subsequent load changes during ongoing operation induce smaller error fluctuations, as reflected by reduced peak magnitude at 12s and 24s, which can be attributed to the membrane potential dynamics of spiking neurons that preserve historical information and facilitate smoother adaptation.
-Fig.
-6b, 6c and 6d further show that the convergence process consistently exhibits three distinct error peaks following each load change.
-These peaks directly correspond to the hierarchical loop structure of the framework: the spinal cord module (first-order level) updates at the control frequency and therefore corresponds to every point on the curve; the brainstem and thalamus modules (second-order level) yield the earlier two peaks which reflect the learning convergence process; and the cerebellum module (third-order level) produces the latest peak owing to its longest processing cycle.
-The temporal separation of these peaks validates the layered organization of the control framework.
-To verify this correspondence, ablation experiments were conducted by selectively disabling individual modules, as shown in Fig.
-6e, 6f and 6g.
-Removing the brainstem’s learning capability by fixing feedback control coefficients reduces the number of peaks to two and delays the first peak, indicating diminished adaptability to load change.
-Similarly, disabling the thalamus module by fixing mode-combination weights eliminates its characteristic peak, leaving only brainstem and cerebellar responses.
-Finally, replacing the cerebellum’s adaptive torque generation with fixed compensation suppresses the cerebellar peak and weakens the thalamic one, since the thalamus relies on cerebellar outputs to optimize weight assignment.
-These results confirm the functional contributions of each module and their hierarchical cooperation.
-In summary, the proposed framework exhibits robust adaptability and rapid convergence to load variations, with distinct loop-specific responses emerging in a biologically consistent sequence.
-This step-by-step activation phenomenon mirrors the progression of human neural responses to perturbations, namely M1 reflexes, M2 reflexes, and voluntary movements [forgaard2015voluntary], thereby reinforcing the plausibility and effectiveness of the neuromorphic design.
-V Experiment results
-This section will present real-world experiments on a robotic arm platform to demonstrate contribution (4).
-All spiking neurons, except those in the spinal cord module, are implemented in Python using an open-source SNN learning framework, while the spinal cord module is realized in C++ through direct modeling of membrane potential dynamics.
-The deep learning framework, SNN framework, and computing hardware are identical to that used in the simulation.
+6e, removing the learning capability of the brainstem module by fixing the feedback control coefficients reduces the number of peaks to two and delays the first peak.
+Similarly, as illustrated in Fig.
+6f, disabling the thalamus module by fixing the motor pattern weighting coefficients eliminates the second peak, leaving only the responses associated with the brainstem and cerebellum modules.
+Finally, as shown in Fig.
+6g, replacing the cerebellum network output with a fixed compensation torque suppresses the cerebellum module peak and weakens the thalamus module response, since it relies on cerebellum module outputs to optimize motor pattern integration.
+In summary, these results demonstrate that the proposed framework exhibits strong and progressively improving resilience during operation.
+This capability arises from the synergistic interaction of its hierarchical structure and modular design.
+Removing any module, control level, or pathway leads to a noticeable degradation in performance.
+V Practical results
+This section will present practical experiment results on a commercial robotic arm platform to demonstrate the precision and adaptation performance.
 The robotic arm platform is a Flexiv Rizon 4s equipped with a Flexiv GRAV gripper, as shown in Fig.
 7.
+The deep learning framework, SNN framework, control cycle of each level, and computing hardware are identical to that used in the simulation, and the Flexiv RDK version is 1.5.1.
 Figure 7: The experiment setting.
-The experiments are conducted under nine load conditions, including two material types (sand-filled and water-filled containers) and five mass levels ranging from 1.5 kg (gripper with empty container) to 3.5 kg (container with 2 kg load).
-The upper limit is chosen below the 4 kg load capacity of the robotic platform to maintain a safety margin.
-Eight end-effector trajectories (T1–T8, defined in the Appendix) are tested, all with vertically downward orientation and origin at the base flange center.
-Among them, T1–T5 represent slower, lower-dynamic trajectories, while T6–T8 correspond to faster, higher-dynamic ones.
-For statistical validity, each trajectory–load–controller combination is repeated 10 times.
-In all experiments presented in this section, the trajectory tracking effect is evaluated using the RMSE of the end-effector position and orientation.
+To represent different tasks, five load masses, two load materials, and eight end-effector trajectories are designed.
+The load masses range from 1.5 kg (the gripper with an empty container) to 3.5 kg (the container with a 2 kg load), with increments of 500 g.
+The two load materials are sand and water.
+All trajectories are designed with a vertically downward orientation, with their definition and parameters provided in the Appendix.
+Among them, T1–T5 represent slower trajectories, whereas T6–T8 correspond to faster trajectories.
+The framework is evaluated on each trajectory ten times under different combinations of load mass and material.
+In all experiments presented in this section, the trajectory tracking error is evaluated using the RMSE of the end-effector position and orientation.
 At each time step, the position and orientation RMSE are defined as
@@
 $\sqrt{(\theta_{d}(t)-\theta(t))^{2}+(\psi_{d}(t)-\psi(t))^{2}+(\phi_{d}(t)-\phi(t))^{2}}$
-where $x_{d},y_{d},z_{d},\theta_{d},\psi_{d},\phi_{d}$ denote the desired end-effector position in Cartesian space and the desired pitch, yaw, and roll angles, respectively, and $x,y,z,\theta,\psi,\phi$ denote the corresponding actual values.
+where $x_{d},y_{d},z_{d},\theta_{d},\psi_{d},\phi_{d}$ denote the desired end-effector position in Cartesian space and the desired pitch, yaw, and roll angles, respectively, and $x,y,z,\theta,\psi,\phi$ denote the actual position in Cartesian space and the actual pitch, yaw, and roll angles, respectively.
 The average position and orientation RMSE of a trajectory are defined as
-$\text{RMSE}_{p}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{p}(t)$ (31)
-$\text{RMSE}_{o}=\frac{1}{N}\sum_{t=1}^{N}\text{RMSE}_{o}(t)$ (32)
-where $N$ represents the trajectory length.
-Figure 8: Trajectory tracking result under different trajectory and load conditions.
+$\text{RMSE}_{p}=\frac{1}{L}\sum_{t=1}^{L}\text{RMSE}_{p}(t)$ (31)
+$\text{RMSE}_{o}=\frac{1}{L}\sum_{t=1}^{L}\text{RMSE}_{o}(t)$ (32)
+where $L$ represents the trajectory length.
+Figure 8: Precision and adaptation performance under different trajectory and load settings.
 (a) Position RMSE under sand load.
@@
 Fig.
-8a and 8b show the trajectory tracking errors of the robotic arm when the end-effector container is filled with sand, under various trajectory and load conditions.
-Red and blue bars represent the proposed method and the factory-installed industrial-grade Flexiv Cartesian Position Control, respectively.
-Cross-hatched and slanted-hatched bars indicate the error ranges of trajectories T1–T5 (lower dynamic group) and T6–T8 (higher dynamic group) for a given load, while error bars denote the extrema of the errors.
-Four observations can be drawn:
+8 shows the trajectory tracking errors of the robotic arm under various trajectory and load conditions.
+Fig.
+8a-b correspond to sand loads, while Fig.
+8c-d correspond to water loads.
+The red and blue bars represent the proposed method and the factory-installed Flexiv Cartesian Position Control, respectively.
+Each bar indicates the range of average errors obtained by repeating each trajectory ten times under a given load condition, while the error bars denote the overall range of errors across all trajectories in the corresponding group.
+The slanted-hatched bars correspond to the fast trajectory group (T1–T5), whereas the cross-hatched bars represent the slow trajectory group (T6–T8).
+Five observations can be drawn:
 1.
-Higher tracking accuracy: For any load mass, the proposed method yields smaller mean position and orientation errors, as indicated by the consistently lower red bars compared with the blue bars.
-Quantitatively, the Flexiv method achieves average errors of $1.789\times 10^{-2}$ m (position) and $1.685\times 10^{-2}$ rad (orientation) across all trajectories and load conditions, whereas the proposed method achieves $1.447\times 10^{-2}$ m and $1.473\times 10^{-2}$ rad, corresponding to reductions of 19.1% and 12.6%, respectively.
-These results demonstrate that the proposed method achieves higher control accuracy, thereby exhibiting superiority in the “Precision” indicator of agility.
+Precision: For all load masses, the proposed method consistently yields smaller average position and orientation errors, as indicated by the uniformly lower red bars compared with the blue bars.
+Under sand-load conditions, the Flexiv controller achieves average errors of $1.800\times 10^{-2}$ m and $1.694\times 10^{-2}$ rad across all trajectory and mass configurations, whereas the proposed method reduces these errors to $1.430\times 10^{-2}$ m and $1.489\times 10^{-2}$ rad, corresponding to reductions of 20.6% and 12.1%, respectively.
+Under water-load conditions, the Flexiv controller yields average errors of $1.779\times 10^{-2}$ m and $1.681\times 10^{-2}$ rad, while the proposed method achieves $1.442\times 10^{-2}$ m and $1.519\times 10^{-2}$ rad, resulting in reductions of 18.9% and 9.6%, respectively.
 2.
-Adaptation to load variation: As load mass increases, position errors of the proposed method gradually increase while orientation errors decrease, and both remain generally lower than those of the Flexiv method, as shown by the upward trend of red bars in Fig.
-8a and the downward trend in Fig.
-8b.
-This observation is consistent with the simulation results, confirming that the proposed framework can adapt to arbitrary load masses within a practical range, validating the design of the cerebellar and thalamic modules.
+Adaptation to trajectory shape: Within each trajectory group, the proposed method exhibits greater consistency across different trajectory shapes, as the red bars under both hatching styles are consistently shorter than their blue counterparts.
+For sand-load conditions, the average variation ranges of the mean position and orientation errors for the Flexiv method are $0.441\times 10^{-2}$ m and $0.596\times 10^{-2}$ rad across all mass configurations, whereas those of the proposed method are reduced to $0.146\times 10^{-2}$ m and $0.393\times 10^{-2}$ rad, corresponding to reductions of 66.9% and 34.1%.
+For water-load conditions, the Flexiv method exhibits average variation ranges of $0.455\times 10^{-2}$ m and $0.588\times 10^{-2}$ rad, while the proposed method achieves $0.157\times 10^{-2}$ m and $0.338\times 10^{-2}$ rad, corresponding to reductions of 65.5% and 39.4%, respectively.
 3.
-Consistency across trajectories: Within each trajectory group of similar dynamics, the proposed method exhibits greater consistency across different trajectory shapes, as the red bars for both hatching styles are shorter than their blue counterparts.
-Specifically, the variation ranges of Flexiv errors in position and orientation are $0.447\times 10^{-2}$ m and $0.592\times 10^{-2}$ rad, while those of the proposed method are $0.152\times 10^{-2}$ m and $0.363\times 10^{-2}$ rad, corresponding to reductions of 66.0% and 38.7%.
-This aligns with simulation findings, indicating that the proposed method can rapidly adapt online to arbitrary motion patterns, validating the brainstem and spinal cord modules and demonstrating an advantage in the agility indicator of “Transitions between diverse motion patterns”.
+Adaptation to trajectory speed: The proposed method produces smaller tracking errors for both slow and fast trajectory groups.
+Notably, the improvement is more pronounced for the fast trajectories, as evidenced by the larger gap between the red and blue slanted-hatched bars compared with that between the cross-hatched bars.
+Under sand-load conditions, for the fast trajectory group, the proposed method reduces position and orientation errors by 29.4% and 17.5%, respectively, relative to the Flexiv method—exceeding the average improvements reported in Observation 1.
+Under water-load conditions, corresponding reductions of 28.5% and 14.1% are achieved.
 4.
-Performance under higher dynamics: For the higher-dynamic group, the performance improvement of the proposed method is more pronounced, as the gap between red and blue slanted bars is larger than that between the cross-hatched bars.
-Quantitatively, in the higher-dynamic group, the proposed method reduces position and orientation errors by 24.7% and 17.0% relative to the Flexiv method, exceeding the average improvements reported in conclusion 1.
-This outcome is expected, since the Flexiv method, being model-based, necessarily involves linear simplifications to ensure computational efficiency.
-While such simplifications suffice at low speeds where nonlinear effects (e.g., Coriolis and centrifugal forces) are weak, they deteriorate at higher speeds.
-In contrast, the proposed method adapts online to nonlinear dynamics, leading to smaller error increases.
-These results suggest that the proposed framework is more suitable for fast motions, demonstrating superiority in the agility indicator of “Speed” .
-Fig.
-8c and 8d present results when the container is filled with water.
-Compared to sand, water introduces stronger dynamic variations, thereby exerting greater influence on control performance.
-The proposed method yields mean errors of $1.442\times 10^{-2}$ m (position) and $1.519\times 10^{-2}$ rad (orientation), representing increases of 0.9% and 2.0% relative to the sand condition; the Flexiv method yields $1.779\times 10^{-2}$ m and $1.681\times 10^{-2}$ rad, representing decreases of 1.2% and 0.7%, respectively.
-Both methods show comparable trends across different load materials, indicating the ability to adapt to varying material properties.
-Moreover, the proposed method consistently demonstrates similar control performance under sand and water conditions, and analogous conclusions to the four points above can be drawn, which are omitted here for brevity.
-In summary, the experiment results demonstrate that the proposed framework consistently outperforms the industrial-grade control algorithm across the three key indicators of agility—precision, adaptability to diverse motion patterns, and speed as mentioned in section I—thereby highlighting its effectiveness and the enhanced agility capacity provided by the brain-inspired design.
-VI Conclusion and future works
-In this paper, we propose a CNS-inspired control framework based on SNN, enabling agile robotic arm control in complex environments.
-The framework introduces a hierarchical architecture with five modules implemented using SNNs, and has been validated through both simulations and real-world robotic experiments.
-Results show that this framework operates in a manner similar to the human CNS and can outperform an industrial-grade baseline.
-Future work will focus on improving reinforcement learning for higher reproducibility, enhancing the cerebellar module for better gravity compensation, and further increasing control accuracy.
-[Experiment trajectory setting]
+Adaptation to load mass: As the load mass increases, the position error of the proposed method increases, while the orientation error decreases; nevertheless, both remain consistently lower than those of the Flexiv method.
+This behavior is reflected by the upward trends of the red bars in Fig.
+8a and Fig.
+8c, and the downward trends in Fig.
+8b and Fig.
+8d.
+5.
+Adaptation to load material: Water introduces stronger dynamic variations due to its complex hydrodynamic behavior, thereby exerting a greater influence on control performance.
+Nevertheless, similar trends are observed when comparing Fig.
+8c–d with Fig.
+8a–b, and all of the above observations consistently apply to both load materials.
+Taken together, the experimental results demonstrate that the proposed framework consistently outperforms the factory-installed method across a wide range of task conditions, including variations in trajectory shape, execution speed, load mass, and load material.
+The framework achieves higher precision and stronger adaptation to unstructured changes, with particularly advantages in high-speed scenarios.
+VI Conclusion and future work
+In this paper, we propose a CNS-inspired, SNN-based motion control framework aimed at achieving agile manipulation in unstructured environments.
+The framework adopts a hierarchical architecture comprising five modules, three control levels, and two pathways, and is validated through both simulation studies and practical experiments.
+Taken together, the results demonstrate that the proposed method operates in a manner analogous to the human CNS and fulfills three out of the four key requirements of agile motion control outlined in Section I, namely precision, adaptation, and resilience.
+Energy efficiency is not explicitly analyzed in this paper, as it depends on multiple factors, including the deep learning and SNN frameworks, computational hardware, and the controlled robotic arm platform.
+In addition, orientation tracking performance is consistently inferior to position tracking performance, indicating room for further improvement.
+Moreover, adaptation to varying load masses can be further enhanced, as tracking precision degrades under changes in load mass.
+These aspects, along with a more thorough analysis of velocity tracking error, will be investigated in future work.
+[Practical experiment trajectory setting]
 $\text{T1:}\begin{cases}x=x_{0}+R_{0}\cos(\frac{2\pi t}{T_{0}})\\ y=y_{0}+R_{0}\sin(\frac{2\pi t}{T_{0}})\\ z=z_{0}\end{cases}\text{T2:}\begin{cases}x=x_{1}\\ y=y_{1}+R_{1}\sin(\frac{2\pi t}{T_{1}})\\ z=z_{1}+R_{1}\cos(\frac{2\pi t}{T_{1}})\end{cases}$
@@
 $\theta_{2}$ $\frac{\pi}{6}$ rad $T_{7}$ 3s
-Yanbo Pang received the B.Eng.
-from Sun Yat-sen University in 2022 and the M.Sc.
-from Tsinghua University in 2025.
-He is currently an intern at Booster Robotics.
-His research interests include neuromorphic computing, neurorobotics and robotic arm control.
-Qingkai Li received his B.S.
-from Tsinghua University in 2021.
-He is currently a Ph.D.
-candidate in the Department of Automation at Tsinghua University, with research interests spanning neuromorphic computing, bio-inspired algorithms, imitation learning, and whole-body control.
-Mingguo Zhao received the B.Eng., M.Sc.
-and Ph.D.
-from Harbin Institute of Technology in 1995, 1997 and 2001, respectively.
-He is currently a Full Professor with the Department of Automation and a director of the Beijing Innovation Center for Future Chips, Tsinghua University.
-His research interests include brain-inspired robot control and general robotic platform based on neuromorphic computing and hardware.